### v1

# Imaging Shapes of Atomic Nuclei in High-Energy Nuclear Collisions

Abstract
Atomic nuclei are self-organized, many-body quantum systems bound by strong nuclear forces within femtometer-scale space. These complex systems manifest a diverse set of shapes Bohr and Mottelson (1998); Heyde and Wood (2016), traditionally explored via non-invasive spectroscopic techniques at low energies Cline (1986); Yang et al. (2023). Their instantaneous shapes, obscured by long-timescale quantum fluctuations, are considered not directly observable at low energy. We introduce a complementary method, collective flow assisted nuclear shape imaging, to image the nuclear global shape by colliding them at ultrarelativistic speeds and analyzing the collective response of outgoing debris. This technique captures a collision-specific snapshot of the spatial matter distribution in the nuclei, which, through the hydrodynamic expansion, leaves imprints on the particle momentum distribution patterns observed in detectors Shuryak (2017); Busza et al. (2018). We benchmark this method in collisions of ground state Uranium-238 nuclei, known for its elongated, axial-symmetric shape. Our findings, while confirming an overall deformation broadly consistent with prior low-energy experiments, also indicate a small deviation from axial symmetry in the nuclear ground state. This approach marks a new way of imaging nuclei, especially those with uncertain shape characteristics, and refines initial conditions in high-energy nuclear collisions. It tackles the important issue of nuclear structure evolution across various energy scales.
More than 99.9% of the visible matter in the cosmos resides in the center of atoms – the atomic nuclei composed of nucleons (protons and neutrons). Our present knowledge of their global structure primarily comes from a variety of spectroscopic or scattering experiments Cline (1986); Yang et al. (2023); Hofstadter (1956) at beam energies below hundreds of MeV per nucleon. These studies reveal that most nuclei are ellipsoidally deformed, with greater deformation in nuclei distant from magic numbers (2, 8, 20, 28, 50, 82, and 126) Möller et al. (2016). Investigating this deformation across the nuclear chart has been an important area of research over many decades, and is crucial for topics such as nucleosynthesis Schatz et al. (1998), nuclear fission Schunck and Regnier (2022), and neutrinoless double beta decay ($0\nu\beta\beta$Engel and Menéndez (2017).
Within a collective model picture, the ellipsoidal shape of a nucleus with mass number $A$ is defined in the intrinsic (body-fixed) frame, where its surface $R(\theta,\phi)$ is described by Bohr and Mottelson (1998),
$R(\theta,\phi)$ $=R_{0}(1+\beta_{2}[\cos\gamma Y_{2,0}+\sin\gamma Y_{2,2}])\;.$ (1)
Here, $R_{0}\approx 1.2A^{1/3}$ fm represents the nuclear radius (about 5–8 fm for large nuclei). The spherical harmonics in the real basis $Y_{l,m}(\theta,\phi)$, the quadrupole deformation magnitude $\beta_{2}$, and the triaxiality parameter $\gamma$ define the nuclear shape. The $\gamma$ parameter, spanning $0^{\circ}$ to $60^{\circ}$, controls the ratios of principal radii. Specifically, $\gamma=0^{\circ}$ corresponds to a prolate shape, $\gamma=60^{\circ}$ an oblate shape, and values in between $0^{\circ}<\gamma<60^{\circ}$ to a triaxial shape. While most nuclei are axially-symmetric (prolate or oblate) or have a fluctuating $\gamma$ value ($\gamma$-soft), rigid triaxial shape is uncommon Toh et al. (2013). An example of an axial-symmetric, prolate-deformed nucleus is shown in Fig. 1a.
Nuclear shapes, even in ground states, are not fixed. They exhibit zero-point quantum fluctuations involving various collective and nucleonic degrees of freedom (DOF) at different timescales. These fluctuations superimpose on each other in the laboratory frame. In well-deformed nuclei like ${}^{238}$, dominant fluctuations are in the rotational DOF with a timescale of $\tau_{\mathrm{rot}}\sim I/\hbar\sim 10^{3}$$10^{4}$ fm/$c$ (1 fm/$c=3\times 10^{-24}$ seconds) Nakatsukasa et al. (2016), where $I$ denotes the moment of inertia (Fig. 1b). Consequently, measurement processes in spectroscopic methods, lasting orders of magnitude longer than $\tau_{\mathrm{rot}}$, capture a coherent superposition of wavefunctions in all orientations, obscuring instantaneous nuclear shapes. However, these shapes can be extracted precisely by comparing spectroscopic data (Fig. 1c) with model calculations Heyde and Wood (2011); Bender et al. (2003). Traditional electron-nucleus scattering experiments, though faster than $\tau_{\mathrm{rot}}$, probe mostly localized regions of the nucleus, resulting in an orientation-averaged spherical image after accumulating many events, where the impact of nuclear deformation manifests as a broadening of the charge distribution Hofstadter (1956); Bohr and Mottelson (1998).
Fig. 1: Methods for determining the nuclear shape in low and high energies. a, cartoon of a well-deformed prolate-shaped nucleus. b, quantum fluctuations over Euler angles for this nucleus and associated overall timescale. c, quantum mechanical manifestation of the deformation in terms of the first rotational band of ${}^{238}$. d, aligning the two nuclei in headon body-body configuration (top) and tip-tip configuration (bottom). e, collision of two Lorentz-contracted nuclei at high energy and resulting 3D profile of the initially-produced quark-gluon plasma, where arrows indicate the pressure gradients. f, 3D profile of the quark-gluon plasma at the end of the hydrodynamic expansion before freezes out into particles, where arrows indicate the velocities of fluid cells. g, charged particle tracks measured in the detector. The timescales shown are in units of fm/$c$ – the time for light to travel one femtometer. The body-body configuration has large eccentricity $\varepsilon_{2}$ and small gradient $d_{\perp}$, leading to large elliptic flow $v_{2}$ and smaller average transverse momentum $[p_{\mathrm{T}}]$, and verse versa for tip-tip configuration (see text).
New shape-imaging method. To directly observe the nuclei’s global shape, a measurement must 1) be much quicker than $\tau_{\mathrm{rot}}$ and 2) provide access to the many-body nucleon distribution in each nucleus, which contains the frame-independent shape information. High-energy nuclear collisions, an utterly destructive process, remarkably fulfill these criteria. Conducted at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) with center-of-mass energies per nucleon pair ($\sqrt{s_{\mathrm{NN}}}$) reaching up to 5000 GeV, these collisions completely obliterate the nuclei, temporarily forming a quark-gluon plasma (QGP) – a hot, dense matter of interacting quarks and gluons Riordan and Zajc (2006); Gale et al. (2013); Shuryak (2017); Busza et al. (2018). The nuclear shape influences the QGP’s geometry and its collective expansion, imprinting itself on the momentum distribution of the produced particles. In an ironic twist, this effectively realizes Richard Feynman’s analogy of the seemingly impossible task of “figuring out a pocket watch by smashing two together and observing the flying debris,” and the collective response plays a key role.
Our shape-imaging technique focuses on headon (near zero impact parameter) collisions of prolate-deformed nuclei (Fig. 1d-g). The initial configurations lie between two extremes: body-body (top) and tip-tip (bottom) collisions (Fig. 1d). Prior to impact, Lorentz-contraction flattens the ground state nuclei into pancake-like shapes by a factor of $\Gamma=\frac{1}{2}\mbox{$\sqrt{s_{\mathrm{NN}}}$}/m_{0}>100$, where $m_{0}\approx 0.94$ GeV is the nucleon mass (Fig. 1e). The initial impact, lasting $\tau_{\mathrm{expo}}=2R_{0}/\Gamma\lesssim 0.1$ fm/$c$, acts as an exposure time. The shape and size of the overlap region, reflecting the initially-produced QGP, directly mirror those of the colliding nuclei projected on the transverse ($xy$) plane (Fig. 1e). Body-body collisions create a larger, elongated QGP, which undergoes pressure-gradient-driven expansion (indicated by arrows) until about 10 fm/$c$ Busza et al. (2018), resulting in an inverted, asymmetric distribution (Fig. 1f). In contrast, tip-tip collisions form a compact, circular QGP, driving a more rapid but near symmetric expansion (Fig. 1f). In the final stage, the QGP freezes out into thousands of particles, captured as tracks in detectors, whose angular distributions reflect the initial QGP shape (Fig. 1g). This flow-assisted imaging is akin to the Coulomb Explosion Imaging in molecule structure analysis Vager et al. (1989); Kunitski et al. (2015); Endo et al. (2020); Boll et al. (2022); Miller (2022), where the spatial arrangement of atoms, ionized by an X-ray laser or through passage in thin foils, is deduced from their mutual Coulomb-repulsion-driven expansion. However, the expansion duration in high-energy collisions is about $10^{6}$$10^{9}$ times shorter.
Energy evolution of shapes? A pertinent question arises: How do the nuclear shapes observed in high-energy colliders compare to those derived from low-energy experiments? For well-deformed nuclei like ${}^{238}$, we expect them to align at a basic level. However, there are other correlations (collective vibration, clustering, short-range correlations, etc.) that manifest over increasingly faster timescales from 1000 to a few fm/$c$ Nazarewicz (1998). Moreover, high-energy collisions also probe nuclear structure at subnucleonic level, such as quark and gluon correlations, and modifications caused by dense gluon fields Mäntysaari et al. (2023). As a result, the deformations observed at high energy might differ from those at low energy. These factors are subdominant for well-deformed nuclei but could be more discernible for less-deformed ones. This highlights the need to examine nuclear phenomena across energy scales and discover new phenomena.
Observables. In Fig. 1e, the QGP’s initial shape is quantified by the eccentricity, $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$, calculated from the nucleon distribution in the $xy$-plane, perpendicular to the beam direction. The hydrodynamic expansion, reacting to $\varepsilon_{2}$, results in particle anisotropy, described as $dN/d\phi~{}\propto~{}1+2v_{2}\cos(2\phi)$ aligned with the impact parameter along the $x$-axis. This phenomenon, known as “elliptic flow” ($v_{2}$llitrault (1992), is depicted in Fig. 1g. Additionally, the QGP’s compactness or density gradient, indicated by the overlap’s inverse area $d_{\perp}=1/\sqrt{\left\langle x^{2}\right\rangle\left\langle y^{2}\right\rangle}$ Schenke et al. (2020a), influences the radial expansion or “radial flow”, captured in the event-wise average transverse momentum ($[p_{\mathrm{T}}]$). A key discovery at RHIC was the QGP’s behavior as a nearly perfect, inviscid fluid Adams et al. (2005); Adcox et al. (2005), effectively transforming initial geometry into final state anisotropies. Hydrodynamic models have confirmed linear response relations: $v_{2}\propto\varepsilon_{2}$ Niemi et al. (2013) and $\delta p_{\mathrm{T}}\propto\delta d_{\perp}$ Boźek and Broniowski (2012), where $\delta p_{\mathrm{T}}=[p_{\mathrm{T}}]-\left\langle[p_{\mathrm{T}}]\right\rangle$ and $\delta d_{\perp}=d_{\perp}-\left\langle d_{\perp}\right\rangle$ denote event-wise deviations from mean values.
In collisions with spherical nuclei, non-zero $\varepsilon_{2}$ and $\delta d_{\perp}$ can be generated by the random fluctuations in the position of nucleons in the overlap region. In non-headon collisions, besides these stochastic elements, the overlap region also has an average elliptical shape. This average shape significantly contributes to $\varepsilon_{2}$, known as reaction plane eccentricity $\varepsilon_{2}^{\mathrm{rp}}$ Alver et al. (2008), but has little effect on $\delta d_{\perp}$.
Prolate deformation in the colliding nuclei further modifies $\varepsilon_{2}$ and $d_{\perp}$. Body-body collisions in this context yield high $\varepsilon_{2}$ and low $d_{\perp}$ values, and vice versa for tip-tip collisions. This leads to enhanced, anti-correlated event-by-event fluctuations in $\varepsilon_{2}$ and $d_{\perp}$ Giacalone (2020a), measurable through observables like $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ that are linearly related to the moments of the initial condition $\left\langle\varepsilon_{2}^{2}\right\rangle$, $\left\langle(\delta d_{\perp})^{2}\right\rangle$, and $\left\langle\varepsilon_{2}^{2}\delta d_{\perp}\right\rangle$. These observables, linked to two- and three-body nucleon distributions in the intrinsic frame (Methods), were found to have a simple parametric dependence on the nuclear shape Jia (2022):
$\left\langle v_{2}^{2}\right\rangle$ $=a_{1}+b_{1}\beta_{2}^{2}\;,$
$\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ $=a_{2}+b_{2}\beta_{2}^{2}\;,$
$\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ $=a_{3}-b_{3}\beta_{2}^{3}\cos(3\gamma)\;.$ (2)
The positive coefficients $a_{n}$ and $b_{n}$ reflect the collision geometry and QGP properties. The $b_{n}$ values are nearly independent of the impact parameter, while $a_{n}$ values are minimized in headon collisions, making such collisions ideal for constraining nuclear shape. Our study represents the first quantitative, simultaneous determination of $\beta_{2}$ and $\gamma$ using all three observables in Eq. (2).
[…]
Rigorous calculations of the triaxiality for prolate-like heavy nuclei were limited Hayashi et al. (1984); Sheikh and Hara (1999), with early estimates typically based on simple geometric rotor models, suggesting $\gamma_{\mathrm{U}}=6^{\circ}$$8^{\circ}$ Wu and Cline (1996). A sophisticated energy density functional calculation based on a collective Hamiltonian method (Gogny-5DCH) predicts $\gamma_{\mathrm{U}}\sim 8^{\circ}$ Delaroche et al. (2010). A more recent analysis via a microscopic model suggests a rigid triaxiality with $\gamma_{\mathrm{U}}\sim 15^{\circ}$ Rouoof et al. (2023). However, the origin of triaxiality in these nuclei (rotational or vibrational) remains debated Sharpey-Schafer et al. (2019). Recent advances in the Monte-Carlo Shell Model enabled a more quantitative analysis, indicating a minor rigid triaxial component leading to a small non-zero $\gamma$ for these nuclei, in alignment with spectra data Otsuka et al. (2023). However, triaxiality may change when nuclei are excited, complicating direct spectroscopic verification Otsuka et al. (2023). Thus, measuring ${}^{238}$’s shape could validate our imaging method and directly investigate its ground-state triaxiality.
Fig. 2: Correlation between elliptic flow and radial flow. a, $\left\langle v_{2}^{2}\right\rangle$ versus $\delta p_{\mathrm{T}}/\left\langle[p_{\mathrm{T}}]\right\rangle$ in 0–0.5% most central Au+Au and U+U collisions. b, $\rho_{2}~{}=~{}\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ across centrality, quantifying the strength of $v_{2}$$\delta p_{\mathrm{T}}$ correlation. The elliptic-shaped overlaps in the transverse plane for various centralities are also illustrated.
Our analysis uses U+U data from 2012 and Au+Au data from 2010 and 2011. Each collision produces up to 2000 charged particles in the STAR time-projection chamber (TPC) Anderson et al. (2003), covering the polar angle range $|\theta-90^{\circ}|\lesssim 50^{\circ}$ and full $\phi$ range. The TPC tracks these particles and determines their $p_{\mathrm{T}}$. Collision events are categorized by “centrality”, defined as the percentage of the total inelastic cross-section, with lower percentages indicating more headon collisions. We calculate $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ via established methods Bilandzic et al. (2011) using tracks in $30.2<p_{\mathrm{T}}<$ 3 GeV/$c$. The results incorporate the uncertainties arising from track selection, reconstruction efficiency, background events, and correlations unrelated to flow (Methods).
Results. To directly observe the shape-size correlation depicted in Fig. 1e-f, we analyze the 0–0.5% most central collisions, and correlate $\left\langle v_{2}^{2}\right\rangle$ with event-wise $\delta p_{\mathrm{T}}$ values (Fig. 2a). A pronounced anticorrelation in U+U collisions aligns with the expectation Giacalone (2020a): events with small (large) $\delta p_{\mathrm{T}}$ are enriched with body-body (tip-tip) collisions. This effect is striking, as $\left\langle v_{2}^{2}\right\rangle$ in U+U is twice that of Au+Au at the lowest $\delta p_{\mathrm{T}}$, yet similar at the highest $\delta p_{\mathrm{T}}$.
We quantify this correlation using normalized covariance $\rho_{2}~{}=\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ (Fig. 2b). In Au+Au collisions, $\rho_{2}$ stays relatively constant, with a minor decrease in central regions due to centrality smearing Aad et al. (2023). This smearing can be reduced by averaging over a wider, say 0–5% centrality range Zhou and Jia (2018). In contrast, $\rho_{2}$ in U+U collisions decreases steadily, turning negative at about 7% centrality, reflecting the large prolate deformation of ${}^{238}$. The deformation has the greatest impact in central collisions but also influences other centrality ranges.
Values of observable in one collision system are heavily influenced by QGP properties during hydrodynamic evolution. These final state effects are largely mitigated by taking ratios between the two systems: $R_{\mathcal{O}}=\left\langle\mathcal{O}\right\rangle_{\mathrm{U+U}}/\left\langle\mathcal{O}\right\rangle_{\mathrm{Au+Au}}$ (Methods). Figure 3a–c show ratios for the three observables. Clearly, $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ increase by up to 60% in central collisions, suggesting a large $\beta_{2\rm U}$, while $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ decreases by up to threefold across centralities, indicating both a large $\beta_{2\rm U}$ and a small $\gamma_{\mathrm{U}}$. The ratios in the 0–5% most central range, having the greatest sensitivity to ${}^{238}$ shape, are visualized as hatch bands in Fig. 3d–f.
Fig. 3: Constraining the shape of ${}^{238}$. Left column: ratios of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) between U+U and Au+Au collisions as a function of centrality. The data are compared to the IP-Glasma+MUSIC hydrodynamic model calculation assuming $\beta_{2\rm U}=0.28$ (red) and $\beta_{2\rm U}=0.25$ (blue), whose shaded bands denote the model uncertainties (Methods). Mid-column: ratio values in 0–5% most central collisions (hatch bands) for $\left\langle v_{2}^{2}\right\rangle$ (d) and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (e), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (f) are compared to model calculations as a function of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ for four $\gamma_{\mathrm{U}}$ values. The colored quadrilaterals delineates the allowed ranges of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ from this data-model comparison. g shows the constrained ranges of ($\beta_{2\mathrm{U}}$,$\gamma_{\mathrm{U}}$) from three observables separately, and the confidence contours obtained by combining $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (solid lines). The constraint from $R_{v_{2}^{2}}$ is viewed as a lower limit and hence is not used (see text).
The data are compared with the state-of-the-art IP-Glasma+MUSIC hydrodynamic model Schenke et al. (2020a, b), which combines the fluctuating initial energy density distributions, relativistic viscous hydrodynamics, and hadronic transport. This model, successful in describing various flow observables at both RHIC and the LHC Schenke et al. (2020b), parameterizes nuclear shapes with a deformed Woods-Saxon profile,
$\rho(r,\theta,\phi)\propto[1+\exp(r-R(\theta,\phi))/a]^{-1}.$ (3)
The deformation parameters for Au are fixed to an oblate shape with $\beta_{2\mathrm{Au}}=0.14$, while those for U are varied. The model also considers final state effects by adjusting QGP viscosities and initial condition uncertainties, including variations in nuclear radius $R_{0}$, skin $a$, $\beta_{2\mathrm{Au}}$, and higher-order shapes (Table 1 in Methods). These variations are included in the model uncertainties. The mass numbers of Au and U, differing by only 20%, result in almost complete cancellation of final state effects, leaving model uncertainties predominantly due to variations in initial conditions. In Fig. 3a-c, calculations for $\beta_{2\mathrm{U}}=0.28$ match central data for $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, but overestimates $R_{v_{2}^{2}}$ in 0–30% centrality. This overestimation could stem from the model’s limitations in describing $\varepsilon_{2}^{\mathrm{rp}}$-related $v_{2}$ components, which are more affected by structural parameters like nuclear radius and skin Jia et al. (2023). Thus, the $R_{v_{2}^{2}}$ comparison only sets a lower bound for $\beta_{2\mathrm{U}}$.
In Fig. 3d-f, we contrast the 0–5% central data against predictions for varying $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$. Calculated $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ change linearly with $\beta_{2\mathrm{U}}^{2}$, while $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ follows a $\beta_{2\mathrm{U}}^{3}\cos(3\gamma_{\mathrm{U}})$ trend, aligning remarkably with Eq. (2). The intersections between data and model delineate preferred $\beta_{2\mathrm{U}}$ ranges, yielding $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})=0.240\pm 0.014$ and $\beta_{2\mathrm{U}}(R_{(\delta p_{\mathrm{T}})^{2}})=0.294\pm 0.021$. For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, the favored $\beta_{2\mathrm{U}}$ range varies with $\gamma_{\mathrm{U}}$. These preferred ranges are summarized in Fig. 3g. Since $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})$ value is viewed as a lower limit, a combined analysis of constraints from $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ is performed, leading to $\beta_{2\mathrm{U}}=0.297\pm 0.013$ and $\gamma_{\mathrm{U}}=8.6^{\circ}\pm 4.8^{\circ}$ (mean and one standard deviation, see Methods).
The extracted $\beta_{2\mathrm{U}}$ value is broadly in line with low-energy estimates Raman et al. (1987); Delaroche et al. (2010); Pritychenko et al. (2016), implying other sources of nucleon, quark and gluon correlations in ${}^{238}$ are less impactful compared to its large deformation, as supported by recent model studies Mäntysaari et al. (2023). Meanwhile, the nonzero $\gamma_{\mathrm{U}}$ value, highlighted by the confidence contours in Fig. 3g, suggests a small rigid triaxial component in ${}^{238}$ deformation, consistent with expectations from Ref. Otsuka et al. (2023). This marks the first indication of nuclear ground state triaxiality without involving transitions to excited states.
Applications. We have demonstrated that flow-assisted imaging is a promising tool for exploring the structure of atomic nuclei in their ground state, complementary to the traditional approaches. The method’s strength lies in capturing a fast snapshot of nucleon spatial distribution, applicable to any collision species. This contrasts with nuclear spectroscopy, where complexity and interpretation vary with nucleus’s position on the nuclide chart. This approach is effective for discerning shape differences between species with similar mass numbers, ideally isobar pairs. Many applications are possible, with a few examples given here:
  • Odd-mass nuclides: For odd-mass nuclei, where either $N$ or $Z$ is odd, the nuclear shape is often presumed or determined via laser spectroscopy method Yang et al. (2023) to be similar to adjacent even-even nuclei. However, the spectroscopic data are more complex Bohr and Mottelson (1998) due to the coupling of the single unpaired nucleon with the nuclear core. By comparing the flow observables of odd-mass nuclei to selected even-even neighbors with known shapes, the high-energy approach avoids this complication,
  • Octupole and hexadecapole deformations: These less common and generally weaker deformations Butler and Nazarewicz (1996) can be probed through measurement of higher-order flow harmonics (triangular and quadrangular flows).
  • Dynamic deformations in “soft” nuclei: This method could distinguish between average deformation and transient shape fluctuations via measurements of multi-particle correlations Dimri et al. (2023); Acharya et al. (2021). This technique, sensitive only to the ground state, also sidesteps the complexities of disentangling shape variations during transitions to nuclear excited states.
  • $0\nu\beta\beta$ decay: The decay rate hinges on nuclear matrix elements (NME), significantly affected by the shapes of the initial and final species – a pair of isobars with the same mass number. Present NME uncertainties, partly stemming from inadequate knowledge of nuclear shapes, pose a major challenge in experimental design Engel and Menéndez (2017). This method, tailored for isobars, allows for the precise determination of shape differences between these species. This could reduce NME uncertainties, and hence aid in experiments searching for $0\nu\beta\beta$ decay and enhance our understanding of neutrino properties.
It would be remiss not to mention that this approach also holds promise in advancing the study of QGP, particularly its dynamics and transport properties, which have been limited by a poor understanding of QGP initial conditions Shuryak (2017); Busza et al. (2018) (Fig. 1e). By contrasting flow observables in similarly massed but structurally different species, our technique effectively eliminates final-state effects, thereby isolating initial condition variations seeded by shape differences. This can elucidate mechanisms of initial condition formation, consequently help to improve QGP transport property extraction through Bayesian inferences Bernhard et al. (2019); Everett et al. (2021); Nijs et al. (2021), and lead to breakthroughs in high-energy nuclear physics.
Collective flow assisted nuclear shape imaging is a discovery tool for exploring nuclear structure and high-energy nuclear collision physics. Future research could leverage colliders to conduct experiments with selected isobaric or isobar-like pairs. The combination of high- and low-energy techniques enables interdisciplinary opportunities in the study of atomic nuclei across energy scales.
Methods
.1 Accessing information in the intrinsic frame
The nuclear shape in the intrinsic frame was considered not directly observable in low-energy experiments. In high-energy collisions, the collective flow phenomenon can be used to infer the shape and size of the nucleon distribution in the overlap region of the transverse plane, represented as $\rho\left(\mathbf{r}\right)$ with ${\bf r}=x+iy$, and hence connect directly to the shape characteristics of the two colliding nuclei in their intrinsic frame as discussed below.
The elliptic shape of the heavy-ion initial state is characterized by its amplitude $\varepsilon_{2}$ and direction $\Phi_{2}$, which are defined by nucleon positions as,
$\mathcal{E}_{2}\equiv\varepsilon_{2}e^{2i\Phi_{2}}=\frac{\int_{\mathbf{r}}\mathbf{r}^{2}\rho\left(\mathbf{r}\right)}{\int_{\mathbf{r}}|\mathbf{r}|^{2}\rho(\mathbf{r})},\int_{\mathbf{r}}=\int\mathrm{d}x\mathrm{d}y\;.$ (4)
It is important to note that when the coordinate system is rotated such that $x$ and $y$ coincide with the minor and major axes, the elliptic eccentricity coincides with the usual definition $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$.
The parameter $\varepsilon_{2}$ plays a crucial role in driving the elliptic flow $v_{2}$ of the particle distribution: $v_{2}\propto\varepsilon_{2}$. In general, $n$-particle correlations quantify the $n^{\rm th}$-moments of elliptic flow and radial flow distributions, which are linearly related to the $n^{\rm th}$-moments of $\varepsilon_{2}$ and $d_{\perp}$.
To demonstrate this point, let’s consider collisions at zero impact parameter without loss of generality, where average elliptic geometry vanishes $\left\langle\mathcal{E}_{2}\right\rangle=0$. The second moment of eccentricity for many events becomes Giacalone et al. (2019); Giacalone (2023),
$\left\langle\varepsilon_{2}^{2}\right\rangle=\left\langle\mathcal{E}_{2}\mathcal{E}_{2}^{*}\right\rangle\approx\frac{\int_{\mathbf{r}_{1},\mathbf{r}_{2}}\left(\mathbf{r}_{1}\right)^{2}\left(\mathbf{r}_{2}^{*}\right)^{2}\rho\left(\mathbf{r}_{1},\mathbf{r}_{2}\right)}{\left(\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle\right)^{2}}\;,$ (5)
where $\left\langle\rho(\mathbf{r})\right\rangle$ is the event-averaged profile, and
$\rho\left(\mathbf{r}_{1},\mathbf{r}_{2}\right)=\left\langle\delta\rho(\mathbf{r}_{1})\delta\rho(\mathbf{r}_{2})\right\rangle=\left\langle\rho(\mathbf{r}_{1})\rho(\mathbf{r}_{2})\right\rangle-\left\langle\rho(\mathbf{r}_{1})\right\rangle\left\langle\rho(\mathbf{r}_{2})\right\rangle$
is the usual two-body distribution. Similarly, the third central moments are related to the three-body distribution $\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)=\left\langle\delta\rho(\mathbf{r}_{1})\delta\rho(\mathbf{r}_{2})\delta\rho(\mathbf{r}_{3})\right\rangle$. For example,
$\left\langle\varepsilon_{2}^{2}\delta d_{\perp}/d_{\perp}\right\rangle$ $\approx-\frac{\int_{\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}}\left(\mathbf{r}_{1}\right)^{2}\left(\mathbf{r}_{2}^{*}\right)^{2}|\mathbf{r}_{3}^{2}|\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)}{\left(\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle\right)^{2}\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle}\;,$ (6)
where we denote $\delta d_{\perp}/d_{\perp}\equiv(d_{\perp}-\left\langle d_{\perp}\right\rangle)/\left\langle d_{\perp}\right\rangle$, and the relation $\frac{\delta d_{\perp}}{d_{\perp}}\approx-\frac{\delta\left\langle|\mathbf{r}^{2}|\right\rangle}{\left\langle|\mathbf{r}^{2}|\right\rangle}=-\frac{\int_{\mathbf{r}}|\mathbf{r}^{2}|\delta\rho\left(\mathbf{r}\right)}{\int_{\mathbf{r}}|\mathbf{r}|^{2}\langle\rho(\mathbf{r})\rangle}$ is used.
The quantities $\mathcal{E}_{2}$ and $\delta d_{\perp}/d_{\perp}$ depend not only on the nuclear shape but also the random orientations of the projectile and target nuclei, denoted by Euler angles $\Omega_{p}$ and $\Omega_{t}$. For small quadrupole deformation, it suffices to consider the leading-order forms Jia (2022):
$\frac{\delta d_{\perp}}{d_{\perp}}$ $\approx\delta_{d}+p_{0}(\Omega_{p},\gamma_{p})\beta_{2p}+p_{0}(\Omega_{t},\gamma_{t})\beta_{2t}\;,$
$\mathcal{E}_{2}$ $\approx\mathcal{E}_{0}+{\bm{p}}_{2}(\Omega_{p},\gamma_{p})\beta_{2p}+{\bm{p}}_{2}(\Omega_{t},\gamma_{t})\beta_{2t}.$ (7)
Here the scalar $\delta_{d}$ and vector $\mathcal{E}_{0}\equiv\varepsilon_{0}e^{2i\Phi_{2;0}}$ represent values for spherical nuclei. The values of scalar $p_{0}$ and vector ${\bm{p}}_{2}$ are directly connected to the $xy$-projected one-body distribution $\rho(\mathbf{r})$. Therefore, they depend on the orientation of the two nuclei. The fluctuations of $\delta_{d}$ ($\mathcal{E}_{0}$) are uncorrelated with $p_{0}$ (${\bm{p}}_{2}$). After averaging over collisions with different Euler angles and setting $\beta_{2p}=\beta_{2t}$ and $\gamma_{p}=\gamma_{t}$, we obtain:
$\left\langle\varepsilon_{2}^{2}\right\rangle$ $=\left\langle\varepsilon_{0}^{2}\right\rangle+2\left\langle{\bm{p}}_{2}(\gamma){\bm{p}}_{2}^{*}(\gamma)\right\rangle\beta_{2}^{2}$
$\left\langle(\delta d_{\perp}/d_{\perp})^{2}\right\rangle$ $=\left\langle\delta_{d}^{2}\right\rangle+2\left\langle p_{0}(\gamma)^{2}\right\rangle\beta_{2}^{2}$
$\left\langle\varepsilon_{2}^{2}\delta d_{\perp}/d_{\perp}\right\rangle$ $=\left\langle\varepsilon_{0}^{2}\delta_{d}\right\rangle+2\left\langle p_{0}(\gamma){\bm{p}}_{2}(\gamma){\bm{p}}_{2}(\gamma)^{*}\right\rangle\beta_{2}^{3}\;.$ (8)
It is found that $\left\langle{\bm{p}}_{2}(\gamma){\bm{p}}_{2}^{*}(\gamma)\right\rangle$ and $\left\langle p_{0}(\gamma)^{2}\right\rangle$ are independent of $\gamma$, while $\left\langle p_{0}(\gamma){\bm{p}}_{2}(\gamma){\bm{p}}_{2}(\gamma)^{*}\right\rangle\propto-\cos(3\gamma)$, resulting in expressions in Eq. (2).
The event-averaged moments in Eq. (8) are rotationally invariant and capture the intrinsic many-body distributions of $\rho\left(\mathbf{r}\right)$. Note that the coefficients $a_{n}$ in Eq. (2) are strong functions of centrality that decrease towards central collisions, while coefficients $b_{n}$ change weakly with centrality. Therefore, the impact of deformation is always largest in the most central collisions. In general, it can be shown that the $n$-particle correlations reflect the rotational invariant ${n}^{\mathrm{th}}$ central moments of $\rho\left(\mathbf{r}\right)$, which in turn are connected to the ${n}^{\mathrm{th}}$ moments of the nuclear shape in the intrinsic frame.
.2 Previous experimental attempts on nuclear shapes at high energy
The idea that $v_{2}$ can be enhanced by $\beta_{2}$ was recognized early Li (2000); Shuryak (2000); Filip et al. (2009); Shou et al. (2015); Goldschmidt et al. (2015). Studies at RHIC Adamczyk et al. (2015a) and the LHC ALICE Collaboration (2018); Sirunyan et al. (2019); Aad et al. (2020) in ${}^{238}$+${}^{238}$ and ${}^{129}$e+${}^{129}$e collisions indicated $\beta_{2}$’s influence on $v_{2}$. Several theoretical investigations followed this up to assess the extent to which $\beta_{2}$ can be constrained by $v_{2}$ alone Giacalone et al. (2018); Giacalone (2020b); Giacalone et al. (2021); Ryssens et al. (2023). A challenge with $v_{2}$ is that its $a_{2}$ term in Eq. (2) is impacted by $\varepsilon_{2}^{\mathrm{rp}}$ and often exceeds the $b_{1}\beta_{2}^{2}$ even in central collisions. A recent measurement of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ aimed to assess the triaxiality of ${}^{129}$e Aad et al. (2023), but the extraction of $\gamma_{\mathrm{Xe}}$ was hindered by needing prior knowledge of $\beta_{2\mathrm{Xe}}$ and potentially significant fluctuations in $\gamma_{\mathrm{Xe}}$ Robledo et al. (2008); Bally et al. (2022a, b); Dimri et al. (2023). The combination of multiple observables allows for more quantitative extraction of nuclear shape parameters, aided by Eq. (2).
.3 Event selection
In high-energy experiments, the polar angle $\theta$ is usually mapped to the so-called pseudorapidity variable $\eta=-\ln(\tan(\theta/2))$. The STAR TPC polar angle range $|\theta-90^{\circ}|<50^{\circ}$ corresponds to $|\eta|<1$.
The collision events are selected by requiring a coincidence of signals from two vertex position detectors situated on each side of the STAR barrel, covering a pseudorapidity range of $4.94.4<|\eta|<$ 4.9. To increase the statistics for ultra-central collision (UCC) events, a special sample of Au+Au data in 2010 and U+U data is chosen based on the criteria of high multiplicity in the STAR TPC and minimal activity in the zero-degree calorimeters that cover the beam rapidity Bieser et al. (2003).
During the offline analysis, events are selected to have collision vertices $z_{\mathrm{vtx}}$ within 30 cm of the TPC center along the beam line, and within 2 cm of the beam spot in the transverse plane. Additionally, a selection criterion based on the correlation between the number of TPC tracks and the number of tracks matched to the time-of-flight detector covering $|\eta|<0.9$ is applied to suppress pileup events (events containing more than one collision in the TPC) Llope (2012) and background events.
After applying these selection criteria, the Au+Au data set consists of approximately 528 million minimum-bias events (including 370 million from 2011) and 120 million UCC events. The U+U data set comprises around 300 million minimum-bias events and 5 million UCC events.
.4 Track selection
The charged particle tracks are reconstructed within $|\eta|<1$ using the TPC Anderson et al. (2003). For this analysis, tracks are selected with $|\eta|<1$ and the transverse momentum range $3.00.2<p_{\mathrm{T}}<$ 3.0 GeV/$c$. To ensure good quality, the selected tracks are required to have at least 16 fit points out of a maximum of 45, and the ratio of the number of fit points to the number of possible points must be greater than 0.52. Additionally, to reduce contributions from secondary decays, the distance of the track’s closest approach (DCA) to the primary collision vertex is required to be less than 3 cm.
The tracking efficiency in the TPC was evaluated using the standard STAR Monte Carlo embedding technique Adamczyk et al. (2015b). The efficiencies are nearly independent of $p_{\mathrm{T}}$ for $p_{\mathrm{T}}>0.5$ GeV/$c$, with plateau values ranging from 0.72 (0.69) in the most central Au+Au (U+U) collisions to 0.92 in the most peripheral collisions. The efficiency exhibits some $p_{\mathrm{T}}$-dependent variation, on the order of 10% of the plateau values, within the range of $0.50.2<p_{\mathrm{T}}<$ 0.5 GeV/$c$.
.5 Centrality
The centrality of each collision is determined using $N_{\mathrm{ch}}^{\mathrm{rec}}$, which represents the number of raw reconstructed tracks in $|\eta|<0.5$, satisfying $p_{\mathrm{T}}>0.15$ GeV/$c$, and having more than 10 fit points. After applying a correction to account for the dependence on the collision vertex position and the luminosity, the distribution of $N_{\mathrm{ch}}^{\mathrm{rec}}$ is compared to a Monte Carlo Glauber calculation Adamczyk et al. (2015b). This comparison allows for the determination of centrality intervals expressed as a percentage of the total nucleus-nucleus inelastic cross-section.
.6 Calculation of observables
The $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are calculated using charged tracks as follows:
[…]
$\left\langle v_{2}^{2}\right\rangle$ $=\left\langle\frac{\sum_{i\neq j}w_{i}w_{j}\cos 2(\phi_{i}-\phi_{j})}{\sum_{i\neq j}w_{i}w_{j}}\right\rangle_{\mathrm{evt}}$
[…]
The averages are performed first on all multiplets within a single event and then over all events in a fixed $N_{\mathrm{ch}}^{\mathrm{rec}}$ bin. The track-wise weights $w_{i,j,k}$ account for tracking efficiency and its $\eta$ and $\phi$-dependent variation. The values of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are obtained using the standard method, where particle $i$ and $j$ are selected from the $|\eta|<1$, as well as the two-subevent method, where particle $i$ and $j$ are selected from pseudorapidity ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, respectively.
The covariance $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ is calculated by averaging over all triplets labeled by particle indices $i$, $j$, and $k$. The standard cumulant framework is used to obtain the results instead of directly calculating all triplets Bilandzic et al. (2011). We also calculated $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ using the two- and three-subevent methods Aad et al. (2023). In the two-subevent method, particles $i$ and $j$ are taken from ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, while particle $k$ can be taken from either subevents. In the three-subevent method, all three particles $i$, $j$ and $k$ are taken from distinct $\eta$ ranges, namely $-1<\eta_{i}<-0.4$, $10.4<\eta_{j}<$ 1, and $|\eta_{k}|<0.3$. The inclusion of a pseudorapidity gap between the particle pairs or triplets suppresses the short-range “non-flow” correlations arising from resonance decays and jets Jia et al. (2017).
The calculation of $\rho_{2}~{}=\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ relies on the input values of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. These components and $\rho_{2}$ are shown in Fig. 4 as a function of centrality. In the central region, enhancements of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are observed in U+U collisions relative to Au+Au collisions, which is consistent with the influence of large $\beta_{2\rm U}$. In contrast, the values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are significantly suppressed in U+U compared to Au+Au collisions across the entire centrality range shown. This suppression is consistent with the negative contribution expected for strong prolate deformation of U as described in Eq. (2).
.7 Influence of non-flow correlations
The results in Fig. 4 are obtained using the two-subevent method. However, as discussed earlier, $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ has contributions from non-flow correlations, which can be explored by comparing results obtained from the standard, two-subevent and three-subevent methods. The results for $\rho_{2}$ are shown in Fig. 5. In this figure, $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are calculated using the two-subevent method, so any variations in $\rho_{2}$ arise only from $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
Differences are observed between the standard and subevent methods, particularly in peripheral collisions. The majority of the variations arise from the comparison between the standard and two-subevent methods, while the differences between the two- and three-subevent methods are relatively small. It is worth mentioning that some of these differences can be attributed to longitudinal fluctuations in $[p_{\mathrm{T}}]$ and $v_{2}$ Khachatryan et al. (2015); Aad et al. (2021, 2023).
.8 Systematic uncertainties
Systematic uncertainties are estimated by varying the track quality selections, the $z_{\mathrm{vtx}}$ cuts, examining the influence of pileup, comparing results from periods with different detector conditions, the possible impact of non-flow contributions, and closure test. The influence of track selection criteria is studied by varying the number of fit hits on the track from a minimum of 16 to 19 and by varying DCA cut from $<3$ cm to $<2.5$ cm, resulting in variations of 1–5% for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$. The impacts on $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are up to $2.5\%$ and $4\%$, respectively.
The influence of track reconstruction on the collision vertex is examined by comparing the results for different $|z_{\mathrm{vtx}}|$ cuts, with differences found to be $0.5$–3% for all observables. Comparisons between different data-taking periods, particularly normal and reverse magnetic field runs in Au+Au collisions, show consistency within their statistical uncertainties. The influence of pileup and background events is studied by varying the cut on the correlation between $N_{\mathrm{ch}}^{\mathrm{rec}}$ and the number of hits in the TOF. The influence is found to be 1–3% for $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and reaches 2–10% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. Comparisons are also made between the 2010 and 2011 Au+Au datasets, which have different active acceptances in the TPC. The results are largely consistent with the quoted uncertainties, although some differences are observed, particularly in the central region, where variations reach 5–10% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
A closure test was conducted, wherein the reconstruction efficiency and its variations in $\eta$ and $\phi$ from the data were utilized to retain a fraction of the particles generated from a multi-phase transport model Lin et al. (2005). Subsequently, a track-by-track weight, as described in Eq. 9, was applied to the accepted particles. All observables are calculated using the accepted particles and compared to those obtained using the original particles. This procedure allowed us to recover $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ within their statistical uncertainties. However, a 2–3% nonclosure was observed in $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$. Nevertheless, it’s important to note that such non-closures largely cancel out when considering the ratios between U+U and Au+Au collisions.
Several additional cross-checks were also carried out. The track reconstruction efficiency has an approximate 5% uncertainty due to its reliance on particle type and occupancy dependence. We repeated the analysis by varying this efficiency, and the variations in the results were either less than 1% or consistent with their statistical uncertainties. The reconstructed $p_{\mathrm{T}}$ can differ from the true value due to finite momentum resolution. This effect was investigated by further smearing the reconstructed $p_{\mathrm{T}}$ according to the known resolution, calculating the observable, and comparing the results with the original ones. A discrepancy of approximately 0.5% was observed for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, while other observables remained consistent within their statistical uncertainties. These effects cancel in the ratios between U+U and Au+Au collisions.
Modest differences are observed between the standard method and subvent methods for all observables, particularly towards the peripheral collisions. Some of these differences are attributable to non-flow contributions, while the remainder could arise from longitudinal flow decorrelations. We obtain default results as the average of the standard method and two-subevent method, and half of the differences are assigned as systematic uncertainties. They are 1–2% for $\left\langle v_{2}^{2}\right\rangle$, 1–3% for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and 2–4% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in the 0–40% centrality range, respectively.
The total systematic uncertainties are combined in quadrature from different sources, and in some cases, they are larger than the statistical uncertainties. The total uncertainties of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are 2.5–4%, 2–5%, and 4–10%, respectively. Note that for the ratios between U+U and Au+Au, their uncertainties are evaluated for each source and combined in quadrature to form the total systematic uncertainties. This process results in a partial cancellation of the uncertainties between the two systems.
.9 Assigning uncertainties on $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$
A standard pseudo-experiment procedure, similar to that described in Ref. Nisius (2014), is employed to combine the uncertainties from $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ shown in Fig. 3g. We assume that the total uncertainties extracted from the two observables are independent, and we model the probability density function as follows:
$P(\beta_{2\mathrm{U}},\gamma_{\mathrm{U}})\propto\exp\left(-\frac{(\beta_{2\mathrm{U}}-\bar{\beta}_{a})^{2}}{2\sigma_{a}^{2}}-\frac{(\beta_{2\mathrm{U}}-\bar{\beta}_{b}(\gamma_{\mathrm{U}}))^{2}}{2\sigma_{b}^{2}(\gamma_{\mathrm{U}})}\right)\;.$ (10)
Here, $\bar{\beta}_{a}=0.294$ and $\sigma_{a}=0.021$ represent the mean and uncertainty of $\beta_{2\mathrm{U}}$ extracted from $R_{(\delta p_{\mathrm{T}})^{2}}$ in Fig. 3g. Similarly, $\bar{\beta}_{b}$ and $\sigma_{b}$ are the mean and uncertainty of $\beta_{2\mathrm{U}}$ from $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, and they depend on the parameter $\gamma_{\mathrm{U}}$. We sample a uniform prior distribution in $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$ to obtain the posterior distribution. From this posterior distribution, we obtained the mean and one standard deviation uncertainty of $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$, as well as the confidence contours displayed in Fig. 3g.
We also performed an extraction using all three ratio observables. In this case, an additional Gaussian probability density function derived for $R_{v_{2}^{2}}$ with a mean $\bar{\beta}_{c}=0.240$ and variance $\sigma_{c}=0.014$ is multiplied to Eq. (10), and the same procedure described above is repeated. This gives $\beta_{2\mathrm{U}}=0.281\pm 0.010$ and $\gamma_{\mathrm{U}}=4.4^{\circ}\pm 3.2^{\circ}$. However, we do not use this estimate, given the imperfect description of $v_{2}$ in the hydrodynamic model.
.10 Hydrodynamic model setup and simulation
Table 1 details the Woods-Saxon parameters for Au and U used in the IP-Glasma+MUSIC model calculations. The nucleon-nucleon inelastic cross-sections are chosen to be the standard values 42 mb and 40.6 mb for Au+Au collisions at 200 GeV and U+U collisions at 193 GeV, respectively. For U, the nuclear shape in Eq. (1) is extended to include a possible small axial hexadecapole deformation $\beta_{4}$:
$R(\theta,\phi)$ $=R_{0}(1+\beta_{2}[\cos\gamma Y_{2,0}+\sin\gamma Y_{2,2}]+\beta_{4}Y_{4,0})\;.$ (11)
Most low-energy nuclear structure models favor a modest oblate deformation for ${}^{197}$u Bally et al. (2023) in the range of $\beta_{2\mathrm{Au}}\approx 0.12-0.14$. We assume ${}^{197}$u is oblate shape with $\beta_{2\mathrm{Au}}=0.14$ as the default choice for simulation. For ${}^{238}$, we have scanned several $\beta_{2\mathrm{U}}$ values ranging from 0 to 0.34. We also vary $\beta_{4\mathrm{U}}$ from 0 to 0.09 and $\gamma_{\mathrm{U}}$ in the range of ${}^{\circ}$${}^{\circ}$ to examine the sensitivity of the U+U results to hexadecapole deformation and triaxiality. For each setting, about 100,000–400,000 events are generated using the officially available IP-Glasma+MUSIC Schenke et al. (2020a, b). Each event is oversampled at least 100 times to minimize non-flow effects and statistical fluctuations in the hadronic transport. These calculations were performed using services provided by the Open Science Grid (OSG) Consortium Pordes et al. (2007); Sfiligoi et al. (2009).
The role of final state effects is studied by varying the shear and bulk viscosities simultaneously up and down by 50%. The impact on $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are shown for Au+Au collisions in top of panels of Fig. 6. The values of these flow observables are changed by more than a factor of two as a function of centrality, yet the ratios between U+U and Au+Au collisions, shown in the bottom panels, are relatively stable. A small reduction of $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ are observed in the non-central collisions, when values of viscosities are halved. However, this change is clearly an overestimate since the calculated flow observables greatly overestimate the data. So in the end, half of the variations of the ratios are included in the model uncertainty.
The main theoretical uncertainties arise from variations in nuclear structure parameters. Parameters common between two collision systems, such as the minimum inter-nucleon distance in nuclei $d_{\mathrm{min}}$, are not expected to contribute to the uncertainty significantly. However, other parameters including nuclear radius $R_{0}$, skin $a$, and higher-order deformation $\beta_{4}$ could be different between Au and U, and hence contribute more to the theoretical uncertainty.
Table 1 provides a list of variations of nuclear structure parameters. The impact of these variations on ratios of flow observables is displayed in Fig. 7. The ratios of flow observables are insensitive to these variations in the most central collisions. Clearly, $R_{v_{2}^{2}}$ is particularly sensitive to skin parameter $a$. This is understandable, as $v_{2}$ has a large contribution from the reaction plane flow, which varies strongly with the value of $a$ Jia et al. (2023).
Model uncertainties for the ratios are derived by combining the impact of varying viscosities, together with various sources from Fig. 7 according to the Barlow prescription Barlow (2002). As a consequence, checks that are consistent with the default calculation within their statistical uncertainties do not contribute to the model uncertainties. The combined model uncertainties for one standard deviation are illustrated in Fig. 3.
A cross-check was also carried out for an alternative hydrodynamic code, the Trajectum model Nijs and van der Schee (2023); Giacalone et al. (2023). This model has 20 parameter sets obtained from a Bayesian analysis of the Pb+Pb data at the LHC, but was unfortunately not tuned to the RHIC data. For this calculation, we simply repeat the calculation at RHIC energy and calculate the same observables. While the description of the $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ is reasonable, several parameter sets give negative values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in mid-central collisions, and subsequently not used. The calculation is performed for the remaining 16 parameter sets as a function of centrality, and root mean square variation among these calculations is assigned as the uncertainty.
Figure 8 displays the ratios of flow observables from Trajectum and compares them to IP-Glasma+MUSIC. The results from these two models agree within their uncertainties for $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$, with Trajectum predictions slightly higher in the UCC region. This leads to slightly lower values of $\beta_{2\mathrm{U}}$: $\beta_{2\mathrm{U}}=0.230\pm 0.010$ for $R_{v_{2}^{2}}$ and $\beta_{2\mathrm{U}}=0.273\pm 0.015$ for $R_{(\delta p_{\mathrm{T}})^{2}}$.
For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, however, the Trajectum model tends to under-predict the data. In central collisions, such discrepancy can be improved by adjusting the triaxiality parameter to $\gamma_{\mathrm{U}}=15^{\circ}$. Overall, the comparison of the Trajectum model with data leads to similar constraints on $\beta_{2\mathrm{U}}$ but slightly larger $\gamma_{\mathrm{U}}$.
Table 1: The choices of Woods-Saxon parameters in Eqs. (3) and (11), including deformations, in the IP-Glasma+MUSIC model. The default values are denoted by bold font, while the rest are variations designed to constrain the values of $(\beta_{2\mathrm{U}},\gamma_{\mathrm{U}})$ and derive theoretical uncertainties associated with other structure parameters.
Species R(fm) a (fm) $d_{\mathrm{min}}$(fm) $\beta_{2}$ $\beta_{4}$ $\gamma$ (${}^{\circ}$)
${}^{197}$u 6.62,6.38 0.52 0.9, 0.4 0.14,0.12 0 60
${}^{238}$ 6.81,7.07 0.55 0.9, 0.4 0,0.15,0.22 0.09,0 0, 10, 15
0.495,0.605 0.25,0.28,0.34 20

### v2

@@
 Atomic nuclei are self-organized, many-body quantum systems bound by strong nuclear forces within femtometer-scale space.
-These complex systems manifest a diverse set of shapes Bohr and Mottelson (1998); Heyde and Wood (2016), traditionally explored via non-invasive spectroscopic techniques at low energies Cline (1986); Yang et al.
+These complex systems manifest a variety of shapes Bohr and Mottelson (1998); Heyde and Wood (2016), traditionally explored using non-invasive spectroscopic techniques at low energies Cline (1986); Yang et al.
 (2023).
-Their instantaneous shapes, obscured by long-timescale quantum fluctuations, are considered not directly observable at low energy.
-We introduce a complementary method, collective flow assisted nuclear shape imaging, to image the nuclear global shape by colliding them at ultrarelativistic speeds and analyzing the collective response of outgoing debris.
-This technique captures a collision-specific snapshot of the spatial matter distribution in the nuclei, which, through the hydrodynamic expansion, leaves imprints on the particle momentum distribution patterns observed in detectors Shuryak (2017); Busza et al.
+However, at these energies, their instantaneous shapes are obscured by long-timescale quantum fluctuations, making direct observation challenging.
+Here we introduce the “collective flow assisted nuclear shape imaging” method, which images the nuclear global shape by colliding them at ultrarelativistic speeds and analyzing the collective response of outgoing debris.
+This technique captures a collision-specific snapshot of the spatial matter distribution within the nuclei, which, through the hydrodynamic expansion, imprints patterns on the particle momentum distribution observed in detectors Shuryak (2017); Busza et al.
 (2018).
-We benchmark this method in collisions of ground state Uranium-238 nuclei, known for its elongated, axial-symmetric shape.
-Our findings, while confirming an overall deformation broadly consistent with prior low-energy experiments, also indicate a small deviation from axial symmetry in the nuclear ground state.
-This approach marks a new way of imaging nuclei, especially those with uncertain shape characteristics, and refines initial conditions in high-energy nuclear collisions.
-It tackles the important issue of nuclear structure evolution across various energy scales.
+We benchmark this method in collisions of ground state Uranium-238 nuclei, known for their elongated, axial-symmetric shape.
+Our findings show a large deformation with a slight deviation from axial symmetry in the nuclear ground state, aligning broadly with previous low-energy experiments.
+This approach offers a new method for imaging nuclear shapes, enhances our understanding of the initial conditions in high-energy collisions and addresses the important issue of nuclear structure evolution across energy scales.
 More than 99.9% of the visible matter in the cosmos resides in the center of atoms – the atomic nuclei composed of nucleons (protons and neutrons).
-Our present knowledge of their global structure primarily comes from a variety of spectroscopic or scattering experiments Cline (1986); Yang et al.
+Our knowledge of their global structure primarily comes from spectroscopic or scattering experiments Cline (1986); Yang et al.
 (2023); Hofstadter (1956) at beam energies below hundreds of MeV per nucleon.
-These studies reveal that most nuclei are ellipsoidally deformed, with greater deformation in nuclei distant from magic numbers (2, 8, 20, 28, 50, 82, and 126) Möller et al.
+These studies show that most nuclei are ellipsoidally deformed, with greater deformation in nuclei distant from magic numbers (2, 8, 20, 28, 50, 82, and 126) Möller et al.
 (2016).
-Investigating this deformation across the nuclear chart has been an important area of research over many decades, and is crucial for topics such as nucleosynthesis Schatz et al.
+Investigating nuclear shape across the Segrè chart has been an important area of research over many decades and is crucial for topics such as nucleosynthesis Schatz et al.
 (1998), nuclear fission Schunck and Regnier (2022), and neutrinoless double beta decay ($0\nu\beta\beta$Engel and Menéndez (2017).
-Within a collective model picture, the ellipsoidal shape of a nucleus with mass number $A$ is defined in the intrinsic (body-fixed) frame, where its surface $R(\theta,\phi)$ is described by Bohr and Mottelson (1998),
+In a collective model picture, the ellipsoidal shape of a nucleus with mass number $A$ is defined in the intrinsic (body-fixed) frame, in which its surface $R(\theta,\phi)$ is described by Bohr and Mottelson (1998),
 $R(\theta,\phi)$ $=R_{0}(1+\beta_{2}[\cos\gamma Y_{2,0}+\sin\gamma Y_{2,2}])\;.$ (1)
-Here, $R_{0}\approx 1.2A^{1/3}$ fm represents the nuclear radius (about 5–8 fm for large nuclei).
+Here $R_{0}\approx 1.2A^{1/3}$ fm represents the nuclear radius.
 The spherical harmonics in the real basis $Y_{l,m}(\theta,\phi)$, the quadrupole deformation magnitude $\beta_{2}$, and the triaxiality parameter $\gamma$ define the nuclear shape.
-The $\gamma$ parameter, spanning $0^{\circ}$ to $60^{\circ}$, controls the ratios of principal radii.
+The $\gamma$ parameter, spanning $0^{\circ}$$60^{\circ}$, controls the ratios of principal radii.
 Specifically, $\gamma=0^{\circ}$ corresponds to a prolate shape, $\gamma=60^{\circ}$ an oblate shape, and values in between $0^{\circ}<\gamma<60^{\circ}$ to a triaxial shape.
-While most nuclei are axially-symmetric (prolate or oblate) or have a fluctuating $\gamma$ value ($\gamma$-soft), rigid triaxial shape is uncommon Toh et al.
+Although most nuclei are axially-symmetric (prolate or oblate) or have a fluctuating $\gamma$ value ($\gamma$-soft), the rigid triaxial shape is uncommon Toh et al.
 (2013).
@@
 These fluctuations superimpose on each other in the laboratory frame.
-In well-deformed nuclei like ${}^{238}$, dominant fluctuations are in the rotational DOF with a timescale of $\tau_{\mathrm{rot}}\sim I/\hbar\sim 10^{3}$$10^{4}$ fm/$c$ (1 fm/$c=3\times 10^{-24}$ seconds) Nakatsukasa et al.
+In well-deformed nuclei like 238U, dominant fluctuations are in the rotational DOF with a timescale of $\tau_{\mathrm{rot}}\sim I/\hbar\sim 10^{3}$$10^{4}$ fm/$c$ (1 fm/$c=3\times 10^{-24}$ seconds $=$ 3 yoctoseconds) Nakatsukasa et al.
 (2016), where $I$ denotes the moment of inertia (Fig.
 1b).
-Consequently, measurement processes in spectroscopic methods, lasting orders of magnitude longer than $\tau_{\mathrm{rot}}$, capture a coherent superposition of wavefunctions in all orientations, obscuring instantaneous nuclear shapes.
-However, these shapes can be extracted precisely by comparing spectroscopic data (Fig.
+Consequently, measurement processes in spectroscopic methods, lasting orders of magnitude longer than $\tau_{\mathrm{rot}}$, capture a coherent superposition of wavefunctions in all orientations.
+Their shapes are usually inferred by comparing spectroscopic data (Fig.
 1c) with model calculations Heyde and Wood (2011); Bender et al.
 (2003).
-Traditional electron-nucleus scattering experiments, though faster than $\tau_{\mathrm{rot}}$, probe mostly localized regions of the nucleus, resulting in an orientation-averaged spherical image after accumulating many events, where the impact of nuclear deformation manifests as a broadening of the charge distribution Hofstadter (1956); Bohr and Mottelson (1998).
+Traditional electron-nucleus scattering experiments, although faster than $\tau_{\mathrm{rot}}$, probe mainly localized regions of the nucleus, giving an orientation-averaged spherical image after accumulating many events, in which the impact of deformation manifests as a broadening of the charge distribution Hofstadter (1956); Bohr and Mottelson (1998).
 Fig.
@@
 b, Quantum fluctuations over Euler angles for this nucleus and associated overall timescale.
-c, quantum mechanical manifestation of the deformation in terms of the first rotational band of ${}^{238}$.
-d, aligning the two nuclei in headon body-body configuration (top) and tip-tip configuration (bottom).
-e, collision of two Lorentz-contracted nuclei at high energy and resulting 3D profile of the initially-produced quark-gluon plasma, where arrows indicate the pressure gradients.
-f, 3D profile of the quark-gluon plasma at the end of the hydrodynamic expansion before freezes out into particles, where arrows indicate the velocities of fluid cells.
+c, Quantum mechanical manifestation of the deformation in terms of the first rotational band of 238U.
+d, Aligning the two nuclei in head-on body-body configuration (top) and tip-tip configuration (bottom).
+e, High-energy collision of two Lorentz-contracted nuclei and resulting 3D profile of the initially-produced quark-gluon plasma, in which the arrows indicate the pressure gradients.
+f, The 3D profile of the quark-gluon plasma at the end of the hydrodynamic expansion before it freezes out into particles, in which arrows indicate the velocities of fluid cells.
 g, Charged particle tracks measured in the detector.
-The timescales shown are in units of fm/$c$ – the time for light to travel one femtometer.
-The body-body configuration has large eccentricity $\varepsilon_{2}$ and small gradient $d_{\perp}$, leading to large elliptic flow $v_{2}$ and smaller average transverse momentum $[p_{\mathrm{T}}]$, and verse versa for tip-tip configuration (see text).
+The timescales shown are in units of fm/$c$ – the time for light to travel 1 femtometer.
+The body-body configuration has large eccentricity $\varepsilon_{2}$ and small gradient $d_{\perp}$, leading to large elliptic flow $v_{2}$ and smaller average transverse momentum $[p_{\mathrm{T}}]$, and vice versa for tip-tip configuration (see main text).
 New shape-imaging method.
-To directly observe the nuclei’s global shape, a measurement must 1) be much quicker than $\tau_{\mathrm{rot}}$ and 2) provide access to the many-body nucleon distribution in each nucleus, which contains the frame-independent shape information.
-High-energy nuclear collisions, an utterly destructive process, remarkably fulfill these criteria.
-Conducted at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) with center-of-mass energies per nucleon pair ($\sqrt{s_{\mathrm{NN}}}$) reaching up to 5000 GeV, these collisions completely obliterate the nuclei, temporarily forming a quark-gluon plasma (QGP) – a hot, dense matter of interacting quarks and gluons Riordan and Zajc (2006); Gale et al.
-(2013); Shuryak (2017); Busza et al.
+To directly observe the global shape of the nuclei, a measurement must 1) be much quicker than $\tau_{\mathrm{rot}}$ and 2) provide access to the many-body nucleon distribution in each nucleus.
+High-energy nuclear collisions, an utterly destructive process, remarkably fulfil these criteria.
+Conducted at the Relativistic Heavy-Ion Collider (RHIC) and the Large Hadron Collider (LHC) with center-of-mass energies per nucleon pair ($\sqrt{s_{\mathrm{NN}}}$) reaching up to 5000 GeV, these collisions completely obliterate the nuclei, temporarily forming a quark-gluon plasma (QGP) – a hot, dense matter of interacting quarks and gluons Shuryak (2017); Busza et al.
 (2018).
-The nuclear shape influences the QGP’s geometry and its collective expansion, imprinting itself on the momentum distribution of the produced particles.
-In an ironic twist, this effectively realizes Richard Feynman’s analogy of the seemingly impossible task of “figuring out a pocket watch by smashing two together and observing the flying debris,” and the collective response plays a key role.
-Our shape-imaging technique focuses on headon (near zero impact parameter) collisions of prolate-deformed nuclei (Fig.
+The nuclear shape influences the geometry of QGP and its collective expansion, imprinting itself on the momentum distribution of the produced particles.
+In an ironic twist, this effectively realizes Richard Feynman’s analogy of the seemingly impossible task of “figuring out a pocket watch by smashing two together and observing the flying debris.” The collective response plays a key role.
+Our shape-imaging technique focuses on head-on (near zero impact parameter) collisions of prolate-deformed nuclei (Fig.
 1d-g).
@@
 1d).
-Prior to impact, Lorentz-contraction flattens the ground state nuclei into pancake-like shapes by a factor of $\Gamma=\frac{1}{2}\mbox{$\sqrt{s_{\mathrm{NN}}}$}/m_{0}>100$, where $m_{0}\approx 0.94$ GeV is the nucleon mass (Fig.
+Before impact, Lorentz-contraction flattens the ground-state nuclei into pancake-like shapes by a factor of $\Gamma=\frac{1}{2}\mbox{$\sqrt{s_{\mathrm{NN}}}$}/m_{0}>100$, where $m_{0}\approx 0.94$ GeV is the nucleon mass (Fig.
 1e).
 The initial impact, lasting $\tau_{\mathrm{expo}}=2R_{0}/\Gamma\lesssim 0.1$ fm/$c$, acts as an exposure time.
-The shape and size of the overlap region, reflecting the initially-produced QGP, directly mirror those of the colliding nuclei projected on the transverse ($xy$) plane (Fig.
+The shape and size of the overlap region, reflecting the initially produced QGP, directly mirror those of the colliding nuclei projected on the transverse ($xy$) plane (Fig.
 1e).
@@
 1f).
-In contrast, tip-tip collisions form a compact, circular QGP, driving a more rapid but near symmetric expansion (Fig.
+By contrast, tip-tip collisions form a compact, circular QGP, driving a more rapid but symmetric expansion (Fig.
 1f).
-In the final stage, the QGP freezes out into thousands of particles, captured as tracks in detectors, whose angular distributions reflect the initial QGP shape (Fig.
+In the final stage, the QGP freezes into thousands of particles, captured as tracks in detectors, whose angular distributions reflect the initial QGP shape (Fig.
 1g).
-This flow-assisted imaging is akin to the Coulomb Explosion Imaging in molecule structure analysis Vager et al.
+This flow-assisted imaging is similar to the Coulomb explosion imaging in molecule structure analysis Vager et al.
 (1989); Kunitski et al.
@@
 (2020); Boll et al.
-(2022); Miller (2022), where the spatial arrangement of atoms, ionized by an X-ray laser or through passage in thin foils, is deduced from their mutual Coulomb-repulsion-driven expansion.
-However, the expansion duration in high-energy collisions is about $10^{6}$$10^{9}$ times shorter.
+(2022); Miller (2022), in which the spatial arrangement of atoms, ionized by an X-ray laser or through passage in thin foils, is deduced from their mutual Coulomb-repulsion-driven expansion.
+However, the expansion duration in high-energy collisions is $10^{6}$$10^{9}$ times shorter.
+The concept that the dynamics of QGP can be utilized to image the geometrical properties of its initial condition was widely recognized.
+This understanding has facilitated the determination of the collision impact parameter and fluctuations in nucleon positions Busza et al.
+(2018), as well as the neutron skin of the colliding nuclei Giacalone et al.
+(2023), by measuring higher-order harmonic flows.
+However, our study took a further step to image the shape of the colliding nuclei through their impact on the initial condition.
 Energy evolution of shapes?
-A pertinent question arises: How do the nuclear shapes observed in high-energy colliders compare to those derived from low-energy experiments?
-For well-deformed nuclei like ${}^{238}$, we expect them to align at a basic level.
-However, there are other correlations (collective vibration, clustering, short-range correlations, etc.) that manifest over increasingly faster timescales from 1000 to a few fm/$c$ Nazarewicz (1998).
-Moreover, high-energy collisions also probe nuclear structure at subnucleonic level, such as quark and gluon correlations, and modifications caused by dense gluon fields Mäntysaari et al.
+A pertinent question is how the shapes observed in high-energy colliders compare with those derived from low-energy experiments.
+For well-deformed nuclei like 238U, we expect them to align at a basic level.
+However, there are other correlations (such as clustering and short-range correlations) that manifest at increasingly faster timescales from 1000 to a few fm/$c$.
+Moreover, high-energy collisions also probe nuclear structure at sub-nucleonic levels, such as quark and gluon correlations, as well as modifications caused by dense gluon fields Mäntysaari et al.
 (2023).
-As a result, the deformations observed at high energy might differ from those at low energy.
-These factors are subdominant for well-deformed nuclei but could be more discernible for less-deformed ones.
-This highlights the need to examine nuclear phenomena across energy scales and discover new phenomena.
+As a result, the deformations observed at high energy may differ from those at low energy, motivating us to examine nuclear phenomena across energy scales and discover new phenomena.
 Observables.
@@
 The hydrodynamic expansion, reacting to $\varepsilon_{2}$, results in particle anisotropy, described as $dN/d\phi~{}\propto~{}1+2v_{2}\cos(2\phi)$ aligned with the impact parameter along the $x$-axis.
-This phenomenon, known as “elliptic flow” ($v_{2}$llitrault (1992), is depicted in Fig.
+This phenomenon, known as “elliptic flow” ($v_{2}$llitrault (1992), is shown in Fig.
 1g.
-Additionally, the QGP’s compactness or density gradient, indicated by the overlap’s inverse area $d_{\perp}=1/\sqrt{\left\langle x^{2}\right\rangle\left\langle y^{2}\right\rangle}$ Schenke et al.
+Moreover, the compactness of the QGP, indicated by the inverse area of the overlap $d_{\perp}\propto 1/\sqrt{\left\langle x^{2}\right\rangle\left\langle y^{2}\right\rangle}$ Schenke et al.
 (2020a), influences the radial expansion or “radial flow”, captured in the event-wise average transverse momentum ($[p_{\mathrm{T}}]$).
-A key discovery at RHIC was the QGP’s behavior as a nearly perfect, inviscid fluid Adams et al.
+A key discovery at RHIC was the behavior of QGP as a nearly perfect, inviscid fluid Adams et al.
 (2005); Adcox et al.
@@
 (2013) and $\delta p_{\mathrm{T}}\propto\delta d_{\perp}$ Boźek and Broniowski (2012), where $\delta p_{\mathrm{T}}=[p_{\mathrm{T}}]-\left\langle[p_{\mathrm{T}}]\right\rangle$ and $\delta d_{\perp}=d_{\perp}-\left\langle d_{\perp}\right\rangle$ denote event-wise deviations from mean values.
-In collisions with spherical nuclei, non-zero $\varepsilon_{2}$ and $\delta d_{\perp}$ can be generated by the random fluctuations in the position of nucleons in the overlap region.
-In non-headon collisions, besides these stochastic elements, the overlap region also has an average elliptical shape.
+In head-on collisions with spherical nuclei, non-zero $\varepsilon_{2}$ and $\delta d_{\perp}$ can be generated by the random fluctuations in the position of nucleons in the overlap region.
+In non-head-on collisions, besides these stochastic elements, the overlap region also has an average elliptical shape.
 This average shape significantly contributes to $\varepsilon_{2}$, known as reaction plane eccentricity $\varepsilon_{2}^{\mathrm{rp}}$ Alver et al.
-(2008), but has little effect on $\delta d_{\perp}$.
-Prolate deformation in the colliding nuclei further modifies $\varepsilon_{2}$ and $d_{\perp}$.
+(2008) but has little effect on the radial quantity $\delta d_{\perp}$.
+Prolate deformation further modifies $\varepsilon_{2}$ and $d_{\perp}$.
 Body-body collisions in this context yield high $\varepsilon_{2}$ and low $d_{\perp}$ values, and vice versa for tip-tip collisions.
-This leads to enhanced, anti-correlated event-by-event fluctuations in $\varepsilon_{2}$ and $d_{\perp}$ Giacalone (2020a), measurable through observables like $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ that are linearly related to the moments of the initial condition $\left\langle\varepsilon_{2}^{2}\right\rangle$, $\left\langle(\delta d_{\perp})^{2}\right\rangle$, and $\left\langle\varepsilon_{2}^{2}\delta d_{\perp}\right\rangle$.
-These observables, linked to two- and three-body nucleon distributions in the intrinsic frame (Methods), were found to have a simple parametric dependence on the nuclear shape Jia (2022):
+This leads to enhanced, anti-correlated event-by-event fluctuations in $\varepsilon_{2}$ and $d_{\perp}$ Giacalone (2020a), measurable through observables such as $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ Bozek (2016) that are linearly related to the moments of the initial condition $\left\langle\varepsilon_{2}^{2}\right\rangle$, $\left\langle(\delta d_{\perp})^{2}\right\rangle$, and $\left\langle\varepsilon_{2}^{2}\delta d_{\perp}\right\rangle$.
+These observables, linked to two- and three-body nucleon distributions in the intrinsic frame (Methods), were found to have a simple parametric dependence on shape parameters Jia (2022):
 $\left\langle v_{2}^{2}\right\rangle$ $=a_{1}+b_{1}\beta_{2}^{2}\;,$
@@
 $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ $=a_{3}-b_{3}\beta_{2}^{3}\cos(3\gamma)\;.$ (2)
-The positive coefficients $a_{n}$ and $b_{n}$ reflect the collision geometry and QGP properties.
-The $b_{n}$ values are nearly independent of the impact parameter, while $a_{n}$ values are minimized in headon collisions, making such collisions ideal for constraining nuclear shape.
-Our study represents the first quantitative, simultaneous determination of $\beta_{2}$ and $\gamma$ using all three observables in Eq.
+The positive coefficients $a_{n}$ and $b_{n}$ capture the collision geometry and QGP properties.
+The $b_{n}$ values are nearly independent of the impact parameter, while $a_{n}$ values are minimized in head-on collisions, making such collisions ideal for constraining nuclear shape.
+Our study offers the first quantitative and simultaneous determination of $\beta_{2}$ and $\gamma$ using all three observables in Eq.
 (2).
-[…]
-Rigorous calculations of the triaxiality for prolate-like heavy nuclei were limited Hayashi et al.
-(1984); Sheikh and Hara (1999), with early estimates typically based on simple geometric rotor models, suggesting $\gamma_{\mathrm{U}}=6^{\circ}$$8^{\circ}$ Wu and Cline (1996).
-A sophisticated energy density functional calculation based on a collective Hamiltonian method (Gogny-5DCH) predicts $\gamma_{\mathrm{U}}\sim 8^{\circ}$ Delaroche et al.
+Nuclear shapes from low energy data.
+Our measurements use data from high-energy 238U+238U and 197Au+197Au collisions.
+These species have contrasting shapes: mildly oblate 197Au (close to magic numbers with $Z$=79 protons and $N$=118 neutrons) and highly prolate 238U (an open shell nucleus with 92 protons and 146 neutrons).
+This comparison helps us to deduce the shape of 238U.
+A state-of-the-art beyond mean-field model, which reproduces the bulk of experimental data on 197Au, predicts deformation values of $\beta_{2\mathrm{Au}}\approx 0.12$–0.14 and $\gamma_{\mathrm{Au}}\approx 43^{\circ}$ Bally et al.
 (2023).
-A more recent analysis via a microscopic model suggests a rigid triaxiality with $\gamma_{\mathrm{U}}\sim 15^{\circ}$ Rouoof et al.
+The deformation of 238U, inferred from measured transition rates within rotational spectra, is estimated at $\beta_{2\mathrm{U}}=0.287\pm 0.007$ Pritychenko et al.
 (2016).
-However, the origin of triaxiality in these nuclei (rotational or vibrational) remains debated Sharpey-Schafer et al.
+Experimental estimates of the Uranium triaxiality have been derived from energy levels and transition data under a rigid-rotor assumption, suggesting $\gamma_{\mathrm{U}}=6^{\circ}$$8^{\circ}$ Wu and Cline (1996).
+One important issue concerns the softness of $\gamma$: whether the nuclei have rigid triaxial shape or fluctuations of $\gamma$ around its mean value Sharpey-Schafer et al.
 (2019).
-Recent advances in the Monte-Carlo Shell Model enabled a more quantitative analysis, indicating a minor rigid triaxial component leading to a small non-zero $\gamma$ for these nuclei, in alignment with spectra data Otsuka et al.
+This issue is complicated by possible changes of $\gamma$ when nuclei are excited Otsuka et al.
 (2023).
-However, triaxiality may change when nuclei are excited, complicating direct spectroscopic verification Otsuka et al.
+Our three-body observable $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in Eq.
+(2) is sensitive only to the mean of the triaxiality, not its fluctuations Dimri et al.
 (2023).
-Thus, measuring ${}^{238}$’s shape could validate our imaging method and directly investigate its ground-state triaxiality.
+Nevertheless, measuring $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$ could validate our imaging method and investigate its ground-state triaxiality.
 Fig.
@@
 b, $\rho_{2}~{}=~{}\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ across centrality, quantifying the strength of $v_{2}$$\delta p_{\mathrm{T}}$ correlation.
-The elliptic-shaped overlaps in the transverse plane for various centralities are also illustrated.
-Our analysis uses U+U data from 2012 and Au+Au data from 2010 and 2011.
-Each collision produces up to 2000 charged particles in the STAR time-projection chamber (TPC) Anderson et al.
-(2003), covering the polar angle range $|\theta-90^{\circ}|\lesssim 50^{\circ}$ and full $\phi$ range.
-The TPC tracks these particles and determines their $p_{\mathrm{T}}$.
-Collision events are categorized by “centrality”, defined as the percentage of the total inelastic cross-section, with lower percentages indicating more headon collisions.
-We calculate $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ via established methods Bilandzic et al.
-(2011) using tracks in $30.2<p_{\mathrm{T}}<$ 3 GeV/$c$.
-The results incorporate the uncertainties arising from track selection, reconstruction efficiency, background events, and correlations unrelated to flow (Methods).
+The elliptic-shaped overlaps in the transverse plane for various centralities are also shown.
+[…]
 Results.
 To directly observe the shape-size correlation depicted in Fig.
-1e-f, we analyze the 0–0.5% most central collisions, and correlate $\left\langle v_{2}^{2}\right\rangle$ with event-wise $\delta p_{\mathrm{T}}$ values (Fig.
+1e–f, we analyze the 0–0.5% most central collisions and correlate $\left\langle v_{2}^{2}\right\rangle$ with event-wise $\delta p_{\mathrm{T}}$ values (Fig.
 2a).
@@
 2b).
-In Au+Au collisions, $\rho_{2}$ stays relatively constant, with a minor decrease in central regions due to centrality smearing Aad et al.
+In Au+Au collisions, $\rho_{2}$ is relatively constant, with a minor decrease in the central region due to centrality smearing Aad et al.
 (2023).
 This smearing can be reduced by averaging over a wider, say 0–5% centrality range Zhou and Jia (2018).
-In contrast, $\rho_{2}$ in U+U collisions decreases steadily, turning negative at about 7% centrality, reflecting the large prolate deformation of ${}^{238}$.
+By contrast, $\rho_{2}$ in U+U collisions decreases steadily, turning negative at about 7% centrality, reflecting the large prolate deformation of 238U.
 The deformation has the greatest impact in central collisions but also influences other centrality ranges.
-Values of observable in one collision system are heavily influenced by QGP properties during hydrodynamic evolution.
-These final state effects are largely mitigated by taking ratios between the two systems: $R_{\mathcal{O}}=\left\langle\mathcal{O}\right\rangle_{\mathrm{U+U}}/\left\langle\mathcal{O}\right\rangle_{\mathrm{Au+Au}}$ (Methods).
+Observables in one collision system are strongly influenced by QGP properties during hydrodynamic evolution.
+By taking ratios between the two systems, such final state effects are largely mitigated: $R_{\mathcal{O}}=\left\langle\mathcal{O}\right\rangle_{\mathrm{U+U}}/\left\langle\mathcal{O}\right\rangle_{\mathrm{Au+Au}}$ (Methods).
 Figure 3a–c show ratios for the three observables.
-Clearly, $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ increase by up to 60% in central collisions, suggesting a large $\beta_{2\rm U}$, while $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ decreases by up to threefold across centralities, indicating both a large $\beta_{2\rm U}$ and a small $\gamma_{\mathrm{U}}$.
-The ratios in the 0–5% most central range, having the greatest sensitivity to ${}^{238}$ shape, are visualized as hatch bands in Fig.
+$R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ increase by up to 60% in central collisions, requiring a large $\beta_{2\rm U}$, whereas $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ decreases by up to threefold across centralities, demanding a large $\beta_{2\rm U}$ and a small $\gamma_{\mathrm{U}}$.
+The ratios in the 0–5% most central range, having the greatest sensitivity to 238U shape, are shown as hatch bands in Fig.
 3d–f.
 Fig.
-3: Constraining the shape of ${}^{238}$.
+3: Constraining the shape of 238U.
 Left column: ratios of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) between U+U and Au+Au collisions as a function of centrality.
-The data are compared to the IP-Glasma+MUSIC hydrodynamic model calculation assuming $\beta_{2\rm U}=0.28$ (red) and $\beta_{2\rm U}=0.25$ (blue), whose shaded bands denote the model uncertainties (Methods).
-Mid-column: ratio values in 0–5% most central collisions (hatch bands) for $\left\langle v_{2}^{2}\right\rangle$ (d) and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (e), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (f) are compared to model calculations as a function of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ for four $\gamma_{\mathrm{U}}$ values.
-The colored quadrilaterals delineates the allowed ranges of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ from this data-model comparison.
-g shows the constrained ranges of ($\beta_{2\mathrm{U}}$,$\gamma_{\mathrm{U}}$) from three observables separately, and the confidence contours obtained by combining $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (solid lines).
-The constraint from $R_{v_{2}^{2}}$ is viewed as a lower limit and hence is not used (see text).
+The data are compared with the IP-Glasma+MUSIC hydrodynamic model calculation assuming $\beta_{2\rm U}=0.28$ (red) and $\beta_{2\rm U}=0.25$ (blue), shaded bands of which denote the model uncertainties (Methods).
+Mid-column: ratio values in 0–5% most central collisions (hatch bands) for $\left\langle v_{2}^{2}\right\rangle$ (d), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (e), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (f) are compared with model calculations as a function of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ for four $\gamma_{\mathrm{U}}$ values.
+The colored quadrilaterals delineate the allowed ranges of $\beta_{2\mathrm{U}}^{2}$ or $\beta_{2\mathrm{U}}^{3}$ from this data-model comparison.
+g, the constrained ranges of ($\beta_{2\mathrm{U}}$,$\gamma_{\mathrm{U}}$) from three observables separately, and the confidence contours obtained by combining $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (solid lines).
+The constraint from $R_{v_{2}^{2}}$ is viewed as a lower limit and hence is not used (see main text).
 The data are compared with the state-of-the-art IP-Glasma+MUSIC hydrodynamic model Schenke et al.
 (2020a, b), which combines the fluctuating initial energy density distributions, relativistic viscous hydrodynamics, and hadronic transport.
-This model, successful in describing various flow observables at both RHIC and the LHC Schenke et al.
+This model, successful in describing flow observables at both RHIC and the LHC Schenke et al.
 (2020b), parameterizes nuclear shapes with a deformed Woods-Saxon profile,
 $\rho(r,\theta,\phi)\propto[1+\exp(r-R(\theta,\phi))/a]^{-1}.$ (3)
-The deformation parameters for Au are fixed to an oblate shape with $\beta_{2\mathrm{Au}}=0.14$, while those for U are varied.
-The model also considers final state effects by adjusting QGP viscosities and initial condition uncertainties, including variations in nuclear radius $R_{0}$, skin $a$, $\beta_{2\mathrm{Au}}$, and higher-order shapes (Table 1 in Methods).
+The parameters for Au are fixed to an oblate shape with $\beta_{2\mathrm{Au}}=0.14$ and $\gamma_{\mathrm{Au}}=45^{\circ}$ Bally et al.
+(2023), whereas those for U are varied.
+The model also considers final state effects by adjusting QGP viscosities and initial condition uncertainties, including variations in nuclear radius $R_{0}$, skin $a$, $\beta_{2\mathrm{Au}}$, $\gamma_{\mathrm{Au}}$, and higher-order shapes (Table 1 in Methods).
 These variations are included in the model uncertainties.
-The mass numbers of Au and U, differing by only 20%, result in almost complete cancellation of final state effects, leaving model uncertainties predominantly due to variations in initial conditions.
+The mass numbers of Au and U, differing by only 20%, result in almost completely canceling final state effects, leaving model uncertainties mainly from initial conditions.
 In Fig.
-3a-c, calculations for $\beta_{2\mathrm{U}}=0.28$ match central data for $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, but overestimates $R_{v_{2}^{2}}$ in 0–30% centrality.
-This overestimation could stem from the model’s limitations in describing $\varepsilon_{2}^{\mathrm{rp}}$-related $v_{2}$ components, which are more affected by structural parameters like nuclear radius and skin Jia et al.
+3a–c, calculations for $\beta_{2\mathrm{U}}=0.28$ match central data for $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, but overestimate $R_{v_{2}^{2}}$ in 0–30% centrality.
+This overestimation stems from the limitations of the model in describing $\varepsilon_{2}^{\mathrm{rp}}$-induced $v_{2}$ components, which are strongly affected by variations in the impact parameter and other structural parameters like nuclear radius and skin Jia et al.
+(2023), as well as possible longitudinal flow decorrelations Jia et al.
 (2024).
-Thus, the $R_{v_{2}^{2}}$ comparison only sets a lower bound for $\beta_{2\mathrm{U}}$.
+Thus, the $R_{v_{2}^{2}}$ comparison is expected to set only a lower bound for $\beta_{2\mathrm{U}}$ in this model.
 In Fig.
 3d-f, we contrast the 0–5% central data against predictions for varying $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$.
-Calculated $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ change linearly with $\beta_{2\mathrm{U}}^{2}$, while $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ follows a $\beta_{2\mathrm{U}}^{3}\cos(3\gamma_{\mathrm{U}})$ trend, aligning remarkably with Eq.
+Calculated $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ change linearly with $\beta_{2\mathrm{U}}^{2}$, whereas $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ follows a $\beta_{2\mathrm{U}}^{3}\cos(3\gamma_{\mathrm{U}})$ trend, aligning remarkably with Eq.
 (2).
-The intersections between data and model delineate preferred $\beta_{2\mathrm{U}}$ ranges, yielding $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})=0.240\pm 0.014$ and $\beta_{2\mathrm{U}}(R_{(\delta p_{\mathrm{T}})^{2}})=0.294\pm 0.021$.
+The intersections between data and model delineate preferred $\beta_{2\mathrm{U}}$ ranges, yielding $\beta_{2\mathrm{U}}(R_{(\delta p_{\mathrm{T}})^{2}})=0.294\pm 0.021$ and a lower limit value $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})=0.234\pm 0.014$.
 For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, the favored $\beta_{2\mathrm{U}}$ range varies with $\gamma_{\mathrm{U}}$.
-These preferred ranges are summarized in Fig.
+These preferred ranges are shown in Fig.
 3g.
-Since $\beta_{2\mathrm{U}}(R_{v_{2}^{2}})$ value is viewed as a lower limit, a combined analysis of constraints from $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ is performed, leading to $\beta_{2\mathrm{U}}=0.297\pm 0.013$ and $\gamma_{\mathrm{U}}=8.6^{\circ}\pm 4.8^{\circ}$ (mean and one standard deviation, see Methods).
-The extracted $\beta_{2\mathrm{U}}$ value is broadly in line with low-energy estimates Raman et al.
-(1987); Delaroche et al.
-(2010); Pritychenko et al.
-(2016), implying other sources of nucleon, quark and gluon correlations in ${}^{238}$ are less impactful compared to its large deformation, as supported by recent model studies Mäntysaari et al.
+A combined analysis of constraints from $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ is performed, yielding $\beta_{2\mathrm{U}}=0.297\pm 0.015$ and $\gamma_{\mathrm{U}}=8.5^{\circ}\pm 4.8^{\circ}$ (mean and one standard deviation, see Methods).
+The data are also compared with Trajectum Giacalone et al.
+(2023), another hydrodynamic model with a different implementation of the initial condition and QGP evolution.
+Trajectum derives constraints on the initial and final state parameters of the QGP based on a Bayesian analysis of the LHC data that are then extrapolated to the RHIC energies.
+These constraints were not tuned to the RHIC data but are nevertheless useful in estimating the theoretical uncertainties.
+The relavent constraints are $\beta_{2\mathrm{U}}=0.275\pm 0.017$ and $\gamma_{\mathrm{U}}=15.5^{\circ}\pm 7.8^{\circ}$.
+A combination of constraints from the two models yields $\beta_{2\mathrm{U}}=0.286\pm 0.025$ and $\gamma_{\mathrm{U}}=8.7^{\circ}\pm 4.5^{\circ}$ (Methods).
+The extracted $\beta_{2\mathrm{U}}$ value is in line with low-energy estimates Pritychenko et al.
+(2016), implying other sources of nucleon, quark and gluon correlations in 238U are less impactful compared with its large deformation, as supported by recent model studies Mäntysaari et al.
 (2023).
-Meanwhile, the nonzero $\gamma_{\mathrm{U}}$ value, highlighted by the confidence contours in Fig.
-3g, suggests a small rigid triaxial component in ${}^{238}$ deformation, consistent with expectations from Ref.
-Otsuka et al.
-(2023).
-This marks the first indication of nuclear ground state triaxiality without involving transitions to excited states.
+Meanwhile, the small $\gamma_{\mathrm{U}}$ value excludes a large triaxiality of Uranium and indicates an average value remarkably consistent with the low energy estimate based on a similar rigid-rotor assumption.
+This marks the first extraction of nuclear ground state triaxiality without involving transitions to excited states.
 Applications.
-We have demonstrated that flow-assisted imaging is a promising tool for exploring the structure of atomic nuclei in their ground state, complementary to the traditional approaches.
-The method’s strength lies in capturing a fast snapshot of nucleon spatial distribution, applicable to any collision species.
-This contrasts with nuclear spectroscopy, where complexity and interpretation vary with nucleus’s position on the nuclide chart.
+The flow-assisted nuclear shape imaging is a promising tool for exploring the structure of atomic nuclei in their ground state.
+The strength of this method lies in capturing a fast snapshot of nucleon spatial distribution, applicable to any collision species.
+This contrasts with nuclear spectroscopy, in which complexity varies with the position of nucleus on the Segrè chart.
 This approach is effective for discerning shape differences between species with similar mass numbers, ideally isobar pairs.
 Many applications are possible, with a few examples given here:
-  • Odd-mass nuclides: For odd-mass nuclei, where either $N$ or $Z$ is odd, the nuclear shape is often presumed or determined via laser spectroscopy method Yang et al.
-(2023) to be similar to adjacent even-even nuclei.
-However, the spectroscopic data are more complex Bohr and Mottelson (1998) due to the coupling of the single unpaired nucleon with the nuclear core.
-By comparing the flow observables of odd-mass nuclei to selected even-even neighbors with known shapes, the high-energy approach avoids this complication,
+  • Odd-mass nuclides: for odd-mass nuclei, where either $N$ or $Z$ is odd, the nuclear shapes should be similar to adjacent even-even nuclei.
+As the transition data are more complex Bohr and Mottelson (1998), the ground-state shapes are usually inferred from the data measured via laser spectroscopy method Yang et al.
+(2023).
+The high-energy approach is suitable for both odd-mass nuclei and even-even nuclei, hence eliminating a possible source of bias in low-energy experiments.
   • Octupole and hexadecapole deformations: these less common and generally weaker deformations Butler and Nazarewicz (1996) can be probed through measurement of higher-order flow harmonics (triangular and quadrangular flows).
   • Dynamic deformations in “soft” nuclei: this method could distinguish between average deformation and transient shape fluctuations via measurements of multi-particle correlations Dimri et al.
-(2023); Acharya et al.
 (2023).
-This technique, sensitive only to the ground state, also sidesteps the complexities of disentangling shape variations during transitions to nuclear excited states.
+For example, the sixth-particle correlator $\left\langle v_{2}^{6}\right\rangle$ has direct sensitivity to the fluctuations of $\gamma$.
+This information was obtained in rare cases at low energy Wu et al.
+(1996); Ayangeakaa et al.
+(2019).
+Our technique, sensitive only to the ground state, also sidesteps the complexities of disentangling shape variations during transitions to excited states.
   • $0\nu\beta\beta$ decay: the decay rate hinges on nuclear matrix elements (NME), significantly affected by the shapes of the initial and final species – a pair of isobars with the same mass number.
-Present NME uncertainties, partly stemming from inadequate knowledge of nuclear shapes, pose a major challenge in experimental design Engel and Menéndez (2017).
-This method, tailored for isobars, allows for the precise determination of shape differences between these species.
+Present NME uncertainties, partly stemming from inadequate knowledge of nuclear shapes, pose a main challenge in experimental design Engel and Menéndez (2017).
+This method, tailored for isobars, allows for precisely determining shape differences between these species.
 This could reduce NME uncertainties, and hence aid in experiments searching for $0\nu\beta\beta$ decay and enhance our understanding of neutrino properties.
-It would be remiss not to mention that this approach also holds promise in advancing the study of QGP, particularly its dynamics and transport properties, which have been limited by a poor understanding of QGP initial conditions Shuryak (2017); Busza et al.
+It would be remiss not to mention that our approach also holds promise in advancing the study of QGP, particularly its dynamics and transport properties, which have been limited by a poor understanding of QGP initial conditions Shuryak (2017); Busza et al.
 (2018) (Fig.
@@
 By contrasting flow observables in similarly massed but structurally different species, our technique effectively eliminates final-state effects, thereby isolating initial condition variations seeded by shape differences.
-This can elucidate mechanisms of initial condition formation, consequently help to improve QGP transport property extraction through Bayesian inferences Bernhard et al.
+This can explain the mechanisms of initial condition formation and consequently help to improve QGP transport property extraction through Bayesian inferences Bernhard et al.
 (2019); Everett et al.
 (2021); Nijs et al.
-(2021), and lead to breakthroughs in high-energy nuclear physics.
+(2021) and lead to breakthroughs in high-energy nuclear physics.
 Collective flow assisted nuclear shape imaging is a discovery tool for exploring nuclear structure and high-energy nuclear collision physics.
 Future research could leverage colliders to conduct experiments with selected isobaric or isobar-like pairs.
-The combination of high- and low-energy techniques enables interdisciplinary opportunities in the study of atomic nuclei across energy scales.
+The combination of high- and low-energy techniques enables interdisciplinary research in the study of atomic nuclei across energy scales.
+.1 Online content
+Any methods, additional references, Nature Portfolio reporting summaries, source data, extended data, supplementary information, acknowledgements, peer review information; details of author contributions and competing interests; and statements of data and code availability are available at https://doi.org/10.1038/s41586-024-08097-2 .
 Methods
-.1 Accessing information in the intrinsic frame
-The nuclear shape in the intrinsic frame was considered not directly observable in low-energy experiments.
-In high-energy collisions, the collective flow phenomenon can be used to infer the shape and size of the nucleon distribution in the overlap region of the transverse plane, represented as $\rho\left(\mathbf{r}\right)$ with ${\bf r}=x+iy$, and hence connect directly to the shape characteristics of the two colliding nuclei in their intrinsic frame as discussed below.
-The elliptic shape of the heavy-ion initial state is characterized by its amplitude $\varepsilon_{2}$ and direction $\Phi_{2}$, which are defined by nucleon positions as,
+.2 Accessing information in the intrinsic frame
+The nuclear shape in the intrinsic frame is not directly observable in low-energy experiments.
+However, in high-energy collisions, the collective flow phenomenon is sensitive to the shape and size of the nucleon distribution in the overlap region of the transverse plane.
+This distribution, denoted as $\rho\left(\mathbf{r}\right)$ with ${\bf r}=x+iy$, provides a direct link to the shape characteristics of the two colliding nuclei in their intrinsic frames, as discussed below.
+The elliptic shape of the heavy-ion initial state is characterized by its amplitude $\varepsilon_{2}$ and direction $\Phi_{2}$, defined by nucleon positions as,
 $\mathcal{E}_{2}\equiv\varepsilon_{2}e^{2i\Phi_{2}}=\frac{\int_{\mathbf{r}}\mathbf{r}^{2}\rho\left(\mathbf{r}\right)}{\int_{\mathbf{r}}|\mathbf{r}|^{2}\rho(\mathbf{r})},\int_{\mathbf{r}}=\int\mathrm{d}x\mathrm{d}y\;.$ (4)
-It is important to note that when the coordinate system is rotated such that $x$ and $y$ coincide with the minor and major axes, the elliptic eccentricity coincides with the usual definition $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$.
-The parameter $\varepsilon_{2}$ plays a crucial role in driving the elliptic flow $v_{2}$ of the particle distribution: $v_{2}\propto\varepsilon_{2}$.
-In general, $n$-particle correlations quantify the $n^{\rm th}$-moments of elliptic flow and radial flow distributions, which are linearly related to the $n^{\rm th}$-moments of $\varepsilon_{2}$ and $d_{\perp}$.
-To demonstrate this point, let’s consider collisions at zero impact parameter without loss of generality, where average elliptic geometry vanishes $\left\langle\mathcal{E}_{2}\right\rangle=0$.
-The second moment of eccentricity for many events becomes Giacalone et al.
+When the coordinate system is rotated such that $x$ and $y$ coincide with the minor and major axes, the elliptic eccentricity coincides with the usual definition $\varepsilon_{2}=\frac{\left\langle y^{2}\right\rangle-\left\langle x^{2}\right\rangle}{\left\langle y^{2}\right\rangle+\left\langle x^{2}\right\rangle}$.
+The parameter $\varepsilon_{2}$ drives the elliptic flow $v_{2}$: $v_{2}\propto\varepsilon_{2}$.
+Let’s now consider collisions at zero impact parameter, where, without loss of generality, the average elliptic geometry vanishes, i.e., $\left\langle\mathcal{E}_{2}\right\rangle=0$.
+The second moment of eccentricity over many events is given by Giacalone et al.
 (2019); Giacalone (2023),
 $\left\langle\varepsilon_{2}^{2}\right\rangle=\left\langle\mathcal{E}_{2}\mathcal{E}_{2}^{*}\right\rangle\approx\frac{\int_{\mathbf{r}_{1},\mathbf{r}_{2}}\left(\mathbf{r}_{1}\right)^{2}\left(\mathbf{r}_{2}^{*}\right)^{2}\rho\left(\mathbf{r}_{1},\mathbf{r}_{2}\right)}{\left(\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle\right)^{2}}\;,$ (5)
-where $\left\langle\rho(\mathbf{r})\right\rangle$ is the event-averaged profile, and
+where $\left\langle\rho(\mathbf{r})\right\rangle$ represents the event-averaged profile, and
 $\rho\left(\mathbf{r}_{1},\mathbf{r}_{2}\right)=\left\langle\delta\rho(\mathbf{r}_{1})\delta\rho(\mathbf{r}_{2})\right\rangle=\left\langle\rho(\mathbf{r}_{1})\rho(\mathbf{r}_{2})\right\rangle-\left\langle\rho(\mathbf{r}_{1})\right\rangle\left\langle\rho(\mathbf{r}_{2})\right\rangle$
 is the usual two-body distribution.
-Similarly, the third central moments are related to the three-body distribution $\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)=\left\langle\delta\rho(\mathbf{r}_{1})\delta\rho(\mathbf{r}_{2})\delta\rho(\mathbf{r}_{3})\right\rangle$.
+Similarly, the third central moments are related to the three-body distribution, $\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)=\left\langle\delta\rho(\mathbf{r}_{1})\delta\rho(\mathbf{r}_{2})\delta\rho(\mathbf{r}_{3})\right\rangle$.
 For example,
-$\left\langle\varepsilon_{2}^{2}\delta d_{\perp}/d_{\perp}\right\rangle$ $\approx-\frac{\int_{\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}}\left(\mathbf{r}_{1}\right)^{2}\left(\mathbf{r}_{2}^{*}\right)^{2}|\mathbf{r}_{3}^{2}|\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)}{\left(\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle\right)^{2}\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle}\;,$ (6)
-where we denote $\delta d_{\perp}/d_{\perp}\equiv(d_{\perp}-\left\langle d_{\perp}\right\rangle)/\left\langle d_{\perp}\right\rangle$, and the relation $\frac{\delta d_{\perp}}{d_{\perp}}\approx-\frac{\delta\left\langle|\mathbf{r}^{2}|\right\rangle}{\left\langle|\mathbf{r}^{2}|\right\rangle}=-\frac{\int_{\mathbf{r}}|\mathbf{r}^{2}|\delta\rho\left(\mathbf{r}\right)}{\int_{\mathbf{r}}|\mathbf{r}|^{2}\langle\rho(\mathbf{r})\rangle}$ is used.
+$\left\langle\varepsilon_{2}^{2}\delta d_{\perp}/d_{\perp}\right\rangle$ $\approx-\frac{\int_{\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}}\left(\mathbf{r}_{1}\right)^{2}\left(\mathbf{r}_{2}^{*}\right)^{2}|\mathbf{r}_{3}^{2}|\rho\left(\mathbf{r}_{1},\mathbf{r}_{2},\mathbf{r}_{3}\right)}{\left(\int_{\mathbf{r}}|\mathbf{r}|^{2}\left\langle\rho(\mathbf{r})\right\rangle\right)^{3}}\;,$ (6)
+where we define $\delta d_{\perp}/d_{\perp}\equiv(d_{\perp}-\left\langle d_{\perp}\right\rangle)/\left\langle d_{\perp}\right\rangle$, and the relation $\frac{\delta d_{\perp}}{d_{\perp}}\approx-\frac{\delta\left\langle|\mathbf{r}^{2}|\right\rangle}{\left\langle|\mathbf{r}^{2}|\right\rangle}=-\frac{\int_{\mathbf{r}}|\mathbf{r}^{2}|\delta\rho\left(\mathbf{r}\right)}{\int_{\mathbf{r}}|\mathbf{r}|^{2}\langle\rho(\mathbf{r})\rangle}$ is used.
 The quantities $\mathcal{E}_{2}$ and $\delta d_{\perp}/d_{\perp}$ depend not only on the nuclear shape but also the random orientations of the projectile and target nuclei, denoted by Euler angles $\Omega_{p}$ and $\Omega_{t}$.
@@
 $\mathcal{E}_{2}$ $\approx\mathcal{E}_{0}+{\bm{p}}_{2}(\Omega_{p},\gamma_{p})\beta_{2p}+{\bm{p}}_{2}(\Omega_{t},\gamma_{t})\beta_{2t}.$ (7)
-Here the scalar $\delta_{d}$ and vector $\mathcal{E}_{0}\equiv\varepsilon_{0}e^{2i\Phi_{2;0}}$ represent values for spherical nuclei.
+Here, the scalar $\delta_{d}$ and vector $\mathcal{E}_{0}$ represent values for spherical nuclei.
 The values of scalar $p_{0}$ and vector ${\bm{p}}_{2}$ are directly connected to the $xy$-projected one-body distribution $\rho(\mathbf{r})$.
@@
 Note that the coefficients $a_{n}$ in Eq.
-(2) are strong functions of centrality that decrease towards central collisions, while coefficients $b_{n}$ change weakly with centrality.
+(2) are strong functions of centrality that decrease towards central collisions, while coefficients $b_{n}$ vary weakly with centrality.
 Therefore, the impact of deformation is always largest in the most central collisions.
 In general, it can be shown that the $n$-particle correlations reflect the rotational invariant ${n}^{\mathrm{th}}$ central moments of $\rho\left(\mathbf{r}\right)$, which in turn are connected to the ${n}^{\mathrm{th}}$ moments of the nuclear shape in the intrinsic frame.
-.2 Previous experimental attempts on nuclear shapes at high energy
+.3 Previous experimental attempts on nuclear shapes at high energy
 The idea that $v_{2}$ can be enhanced by $\beta_{2}$ was recognized early Li (2000); Shuryak (2000); Filip et al.
@@
 (2015).
-Studies at RHIC Adamczyk et al.
-(2015a) and the LHC ALICE Collaboration (2018); Sirunyan et al.
-(2019); Aad et al.
-(2020) in ${}^{238}$+${}^{238}$ and ${}^{129}$e+${}^{129}$e collisions indicated $\beta_{2}$’s influence on $v_{2}$.
-Several theoretical investigations followed this up to assess the extent to which $\beta_{2}$ can be constrained by $v_{2}$ alone Giacalone et al.
+Studies at RHIC Ada (2015) and the LHC ALICE Collaboration (2018); CMS Collaboration (2019); ATLAS Collaboration (2020) in 238U+238U and 129Xe+129Xe collisions indicated $\beta_{2}$’s influence on $v_{2}$.
+Several later theoretical investigations assessed the extent to which $\beta_{2}$ can be constrained by $v_{2}$ alone Giacalone et al.
 (2018); Giacalone (2020b); Giacalone et al.
@@
 (2023).
-A challenge with $v_{2}$ is that its $a_{2}$ term in Eq.
-(2) is impacted by $\varepsilon_{2}^{\mathrm{rp}}$ and often exceeds the $b_{1}\beta_{2}^{2}$ even in central collisions.
-A recent measurement of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ aimed to assess the triaxiality of ${}^{129}$e Aad et al.
-(2023), but the extraction of $\gamma_{\mathrm{Xe}}$ was hindered by needing prior knowledge of $\beta_{2\mathrm{Xe}}$ and potentially significant fluctuations in $\gamma_{\mathrm{Xe}}$ Robledo et al.
+A challenge with $v_{2}$ is that its $a_{1}$ term in Eq.
+(2) is impacted by $\varepsilon_{2}^{\mathrm{rp}}$, which often exceeds the $b_{1}\beta_{2}^{2}$ term even in central collisions.
+A recent measurement of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ aimed to assess the triaxiality of 129Xe Aad et al.
+(2023), but the extraction of $\gamma_{\mathrm{Xe}}$ was hindered by needing prior knowledge of $\beta_{2\mathrm{Xe}}$ and potentially substantial fluctuations in $\gamma_{\mathrm{Xe}}$ Robledo et al.
 (2008); Bally et al.
 (2022a, b); Dimri et al.
+(2023); Fortier et al.
 (2023).
-The combination of multiple observables allows for more quantitative extraction of nuclear shape parameters, aided by Eq.
-(2).
-.3 Event selection
+The combination of several observables in this article allows for more quantitative extraction of nuclear shape parameters.
+.4 Event selection
 In high-energy experiments, the polar angle $\theta$ is usually mapped to the so-called pseudorapidity variable $\eta=-\ln(\tan(\theta/2))$.
@@
 (2003).
-During the offline analysis, events are selected to have collision vertices $z_{\mathrm{vtx}}$ within 30 cm of the TPC center along the beam line, and within 2 cm of the beam spot in the transverse plane.
+In the offline analysis, events are selected to have collision vertices $z_{\mathrm{vtx}}$ within 30 cm of the TPC center along the beamline and within 2 cm of the beam spot in the transverse plane.
 Additionally, a selection criterion based on the correlation between the number of TPC tracks and the number of tracks matched to the time-of-flight detector covering $|\eta|<0.9$ is applied to suppress pileup events (events containing more than one collision in the TPC) Llope (2012) and background events.
-After applying these selection criteria, the Au+Au data set consists of approximately 528 million minimum-bias events (including 370 million from 2011) and 120 million UCC events.
+After applying these selection criteria, the Au+Au dataset has approximately 528 million minimum-bias events (including 370 million in 2011) and 120 million UCC events.
 The U+U dataset comprises around 300 million minimum-bias events and 5 million UCC events.
-.4 Track selection
-The charged particle tracks are reconstructed within $|\eta|<1$ using the TPC Anderson et al.
+.5 Track selection
+For this analysis, tracks are selected with $|\eta|<1$ and the transverse momentum range $3.00.2<p_{\mathrm{T}}<$ 3.0 GeV/$c$.
+To ensure good quality, the selected tracks must have at least 16 fit points out of a maximum of 45, and the ratio of the number of fit points to the number of possible points must be greater than 0.52.
+Additionally, to reduce contributions from secondary decays, the distance of the track’s closest approach (DCA) to the primary collision vertex must be less than 3 cm.
+The tracking efficiency in the TPC was evaluated using the standard STAR Monte Carlo embedding technique Adamczyk et al.
 (2015).
-For this analysis, tracks are selected with $|\eta|<1$ and the transverse momentum range $3.00.2<p_{\mathrm{T}}<$ 3.0 GeV/$c$.
-To ensure good quality, the selected tracks are required to have at least 16 fit points out of a maximum of 45, and the ratio of the number of fit points to the number of possible points must be greater than 0.52.
-Additionally, to reduce contributions from secondary decays, the distance of the track’s closest approach (DCA) to the primary collision vertex is required to be less than 3 cm.
-The tracking efficiency in the TPC was evaluated using the standard STAR Monte Carlo embedding technique Adamczyk et al.
-(2015b).
 The efficiencies are nearly independent of $p_{\mathrm{T}}$ for $p_{\mathrm{T}}>0.5$ GeV/$c$, with plateau values ranging from 0.72 (0.69) in the most central Au+Au (U+U) collisions to 0.92 in the most peripheral collisions.
 The efficiency exhibits some $p_{\mathrm{T}}$-dependent variation, on the order of 10% of the plateau values, within the range of $0.50.2<p_{\mathrm{T}}<$ 0.5 GeV/$c$.
-.5 Centrality
+.6 Centrality
 The centrality of each collision is determined using $N_{\mathrm{ch}}^{\mathrm{rec}}$, which represents the number of raw reconstructed tracks in $|\eta|<0.5$, satisfying $p_{\mathrm{T}}>0.15$ GeV/$c$, and having more than 10 fit points.
-After applying a correction to account for the dependence on the collision vertex position and the luminosity, the distribution of $N_{\mathrm{ch}}^{\mathrm{rec}}$ is compared to a Monte Carlo Glauber calculation Adamczyk et al.
-(2015b).
-This comparison allows for the determination of centrality intervals expressed as a percentage of the total nucleus-nucleus inelastic cross-section.
-.6 Calculation of observables
+After applying a correction to account for the dependence on the collision vertex position and the luminosity, the distribution of $N_{\mathrm{ch}}^{\mathrm{rec}}$ is compared with a Monte Carlo Glauber calculation Adamczyk et al.
+(2015).
+This comparison allows for determining centrality intervals, expressed as a percentage of the total nucleus-nucleus inelastic cross-section.
+.7 Calculation of observables
 The $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are calculated using charged tracks as follows:
@@
 […]
-$\left\langle v_{2}^{2}\right\rangle$ $=\left\langle\frac{\sum_{i\neq j}w_{i}w_{j}\cos 2(\phi_{i}-\phi_{j})}{\sum_{i\neq j}w_{i}w_{j}}\right\rangle_{\mathrm{evt}}$
+$\left\langle v_{2}^{2}\right\rangle$ $=\left\langle\frac{\sum_{i\neq j}w_{i}w_{j}\cos(2(\phi_{i}-\phi_{j}))}{\sum_{i\neq j}w_{i}w_{j}}\right\rangle_{\mathrm{evt}}$
 […]
 The averages are performed first on all multiplets within a single event and then over all events in a fixed $N_{\mathrm{ch}}^{\mathrm{rec}}$ bin.
-The track-wise weights $w_{i,j,k}$ account for tracking efficiency and its $\eta$ and $\phi$-dependent variation.
-The values of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are obtained using the standard method, where particle $i$ and $j$ are selected from the $|\eta|<1$, as well as the two-subevent method, where particle $i$ and $j$ are selected from pseudorapidity ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, respectively.
+The track-wise weights $w_{i,j,k}$ account for tracking efficiency and its $\eta$ and $\phi$ dependent variations.
+The values of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are obtained using the standard method, in which particle $i$ and $j$ are selected from $|\eta|<1$, as well as the two-subevent method, in which particle $i$ and $j$ are selected from pseudorapidity ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, respectively.
+We also calculate the efficiency-corrected charged particle multiplicity in $|\eta|<0.5$, defined as $N_{\mathrm{ch}}=\sum_{i}w_{i}$.
+This observable is used for evaluating the systematics.
 The covariance $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ is calculated by averaging over all triplets labeled by particle indices $i$, $j$, and $k$.
@@
 (2011).
-We also calculated $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ using the two- and three-subevent methods Aad et al.
-(2023).
-In the two-subevent method, particles $i$ and $j$ are taken from ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, while particle $k$ can be taken from either subevents.
-In the three-subevent method, all three particles $i$, $j$ and $k$ are taken from distinct $\eta$ ranges, namely $-1<\eta_{i}<-0.4$, $10.4<\eta_{j}<$ 1, and $|\eta_{k}|<0.3$.
-The inclusion of a pseudorapidity gap between the particle pairs or triplets suppresses the short-range “non-flow” correlations arising from resonance decays and jets Jia et al.
+We also calculated $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ using the two-subevent method Aad et al.
+(2023), where particles $i$ and $j$ are taken from ranges of $-1<\eta_{i}<-0.1$ and $10.1<\eta_{j}<$ 1, while particle $k$ is taken from either subevents.
+Including a pseudorapidity gap between the particle pairs or triplets suppresses the short-range “non-flow” correlations arising from resonance decays and jets Jia et al.
 (2017).
@@
 4 as a function of centrality.
-In the central region, enhancements of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are observed in U+U collisions relative to Au+Au collisions, which is consistent with the influence of large $\beta_{2\rm U}$.
-In contrast, the values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are significantly suppressed in U+U compared to Au+Au collisions across the entire centrality range shown.
+In the central region, enhancements of $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are observed in U+U relative to Au+Au collisions, which is consistent with the influence of large $\beta_{2\rm U}$.
+In contrast, the values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are significantly suppressed in U+U compared with Au+Au collisions across the entire centrality range shown.
 This suppression is consistent with the negative contribution expected for strong prolate deformation of U as described in Eq.
 (2).
-.7 Influence of non-flow correlations
-The results in Fig.
-4 are obtained using the two-subevent method.
-However, as discussed earlier, $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ has contributions from non-flow correlations, which can be explored by comparing results obtained from the standard, two-subevent and three-subevent methods.
-The results for $\rho_{2}$ are shown in Fig.
-5.
-In this figure, $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ are calculated using the two-subevent method, so any variations in $\rho_{2}$ arise only from $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
-Differences are observed between the standard and subevent methods, particularly in peripheral collisions.
-The majority of the variations arise from the comparison between the standard and two-subevent methods, while the differences between the two- and three-subevent methods are relatively small.
-It is worth mentioning that some of these differences can be attributed to longitudinal fluctuations in $[p_{\mathrm{T}}]$ and $v_{2}$ Khachatryan et al.
-(2015); Aad et al.
-(2021, 2023).
-.8 Systematic uncertainties
-Systematic uncertainties are estimated by varying the track quality selections, the $z_{\mathrm{vtx}}$ cuts, examining the influence of pileup, comparing results from periods with different detector conditions, the possible impact of non-flow contributions, and closure test.
+Fig.
+4: Components involved in $v_{2}$-$[p_{\mathrm{T}}]$ correlations.
+The centrality dependences of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) and $\rho_{2}~{}=~{}\frac{\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle}{\left\langle v_{2}^{2}\right\rangle\sqrt{\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle}}$ (d) in U+U (red) and Au+Au (black) collisions.
+They are calculated using the two-subevent method.
+The error bars and shaded bands are statistical and systematic uncertainties, respectively.
+In this analysis, the default results are obtained using the two-subevent method.
+The differences between the standard and two-subevent methods are used to evaluate the impact of non-flow correlations discussed below.
+.8 Influence of non-flow correlations
+An important background in our measurement is non-flow: correlations among a few particles originated from a common source, such as resonance decays and jets, which are uncorrelated with the initial geometry.
+Two approaches are used to estimate the non-flow contributions.
+Non-flow correlations are short-range in $\eta$ and can be suppressed via the subevent method by requiring a rapidity gap between the pairs or triplets of particles in Eq.(9).
+Hence, in the first approach, the differences between the standard and subevent methods provide an estimate of the non-flow contribution.
+However, one must keep in mind that part of the rapidity gap dependence of the signal in central collisions may arise from longitudinal fluctuations in $[p_{\mathrm{T}}]$ and $v_{2}$ due to variations in the initial geometry in $\eta$ Aad et al.
+(2023).
+The second approach assumes that the clusters causing non-flow correlations are mutually independent.
+In this independent-source scenario, non-flow in $n$-particle cumulants is expected to be diluted by the charged particle multiplicity as $1/N_{\mathrm{ch}}^{n-1}$ Borghini et al.
+(2001).
+Therefore, non-flow (nf) contributions in a given centrality can be estimated by,
+$\left\langle v_{2}^{2}\right\rangle_{\mathrm{nf}}\approx\frac{\left[\left\langle v_{2}^{2}\right\rangle N_{\mathrm{ch}}\right]_{\mathrm{peri}}}{N_{\mathrm{ch}}}\;,$
+$\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle_{\mathrm{nf}}\approx\frac{\left[\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle N_{\mathrm{ch}}\right]_{\mathrm{peri}}}{N_{\mathrm{ch}}}\;,$
+$\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle_{\mathrm{nf}}\approx\frac{\left[\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle N_{\mathrm{ch}}^{2}\right]_{\mathrm{peri}}}{N_{\mathrm{ch}}^{2}}$ (10)
+where the subscript “peri” is a label for the peripheral bin.
+This procedure makes two assumptions that are not fully valid: 1) the signal in the peripheral bin is all non-flow, and 2) non-flow in other centralities is unmodified by final state medium effects.
+For example, the medium effects strongly suppress the jet yield and modify the azimuthal structure of non-flow correlations.
+Hence, this approach only provides a qualitative estimate of the non-flow.
+Moreover, this approach is not applicable for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, as medium effects are expected to reduce the momentum differences of non-flow particles, since they are out of local equilibrium.
+Figure 5 shows the $N_{\mathrm{ch}}$-scaled values of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ as a function of centrality in Au+Au collisions.
+The requirement of subevent reduces the signal in the most peripheral bin by 50%, 40%, and 80%, respectively, which can be treated as the amount of non-flow rejected by the subevent requirement.
+Therefore, we use the differences between the standard and subevent methods to estimate the non-flow in the subevent method.
+These differences vary with centrality because of the combined effects of medium modification of non-flow and longitudinal flow decorrelations Aad et al.
+(2021).
+These differences are propagated to the ratios of these observables between U+U and Au+Au.
+They are found to be 1.1%, 3.5%, and 11% for $R_{v_{2}^{2}}$, $R_{(\delta p_{\mathrm{T}})^{2}}$, and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
+Fig.
+5: Impact of non-flow correlations.
+The centrality dependence of $N_{\mathrm{ch}}$ scaled quantities in Au+Au collisions, $\left\langle v_{2}^{2}\right\rangle N_{\mathrm{ch}}$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle N_{\mathrm{ch}}$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle N_{\mathrm{ch}}^{2}$ (c) from the standard method (black solid markers), two-subevent method (red solid markers) and their differences (open markers).
+The error bars are statistical uncertainties.
+The decrease of $\left\langle v_{2}^{2}\right\rangle$-difference towards central collisions is attributed to two track reconstruction effects in the standard method, which does not affect the other two observables.
+The dashed lines indicate the upper limit of the non-flow contribution, assuming $N_{\mathrm{ch}}$-scaling of Eq.
+(10).
+Fig.
+6: Summary of the systematic uncertainties.
+The centrality dependence of ratios $R_{v_{2}^{2}}$ (a), $R_{(\delta p_{\mathrm{T}})^{2}}$ (b), and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ (c) from the two-subevent method with the systematic uncertainties arising from the non-flow estimate (red bars) and other sources (black boxes).
+Figure 5 also provides an estimate of non-flow based on the $N_{\mathrm{ch}}$-scaling method.
+We assume that the entire signals in the 80–100% centrality in two-subevent are non-flow, and then use Eq.
+(10) to estimate the fraction of non-flow as a function of centrality.
+As mentioned earlier, we employ this approach for $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$, where the medium effects may redistribute non-flow correlations in azimuthal angle, instead of suppressing them.
+This approach is unsuitable for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, for which the medium effects should always suppress the non-flow contribution.
+In the 0-5% most central collisions, the estimated non-flow is about 6% for $\left\langle v_{2}^{2}\right\rangle$ and only about 1.4% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
+These differences, when propagated to the ratios, are reduced for $R_{v_{2}^{2}}$, which is positive, and increased for $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, which is negative.
+They amount to about 2.8% for $R_{v_{2}^{2}}$ and 2.5% for $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
+The non-flow systematic uncertainties are taken as the larger of the two approaches for $R_{v_{2}^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, while for $R_{(\delta p_{\mathrm{T}})^{2}}$, the difference between standard and subevent methods is used.
+The total non-flow uncertainties in the 0–5% centrality are 2.8%, 3.5% and 11% for $R_{v_{2}^{2}}$, $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
+Figure 6 contrasts the non-flow systematic uncertainties with other sources of uncertainties (next section) in this analysis.
+In the 0–5% centrality, the non-flow uncertainties are comparable or somewhat larger than other sources, while they are subdominant in other centrality ranges.
+In the literature, non-flow contributions are sometimes estimated using the HIJING model Gyulassy and Wang (1994), which has only non-flow correlations.
+The latter were found to follow very closely Eq.
+(10Bhatta et al.
+(2022); Wang (2024).
+In our second approach, instead of relying on the HIJING model, we assume this $N_{\mathrm{ch}}$-scaling behavior but use real peripheral data as the baseline for non-flow contributions.
+Our findings indicate that the HIJING model fails to quantitatively capture the features of non-flow.
+Specifically, HIJING predicts a much weaker $\Delta\eta$ dependence for $\left\langle v_{2}^{2}\right\rangle$, with only a 13% difference between the standard and two-subevent method, whereas the data indicate a 50% decrease STAR Collaboration (2023) (Fig.
+25 in Ref.
+STAR Collaboration (2023) for $p+p$ collisions).
+Furthermore, we found that the values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ predicted by HIJING are three times larger than the data in peripheral Au+Au collisions.
+Therefore, the non-flow estimation based on the HIJING model in Ref.
+Wang (2024) appears to be exaggerated.
+A more recent estimate Jia et al.
+(2024), based on a transport model incorporating full medium dynamics and Eq.
+(10), yields a non-flow fraction consistent with STAR data.
+This study also suggests a potential increase of the $R_{v_{2}^{2}}$ with $\Delta\eta$ associated with flow decorrelation effects.
+Understanding non-flow correlations as a physical process has always been a work in progress.
+As our knowledge deepens, the non-flow uncertainties are expected to reduce.
+Rather than merely contributing to experimental uncertainties or even being corrected for in the data, non-flow physics should ultimately be incorporated into hydrodynamic models.
+Currently, these models include non-flow effects from resonance decays but lack contributions from jet fragmentation.
+.9 Systematic uncertainties
+Systematic uncertainties include an estimate of the non-flow contributions discussed above and other sources accounting for detector effects and analysis procedure.
+These other sources are estimated by varying the track quality selections, the $z_{\mathrm{vtx}}$ cuts, examining the influence of pileup, comparing results from periods with different detector conditions, and closure test.
 The influence of track selection criteria is studied by varying the number of fit hits on the track from a minimum of 16 to 19 and by varying DCA cut from $<3$ cm to $<2.5$ cm, resulting in variations of 1–5% for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$.
 The impacts on $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are up to $2.5\%$ and $4\%$, respectively.
-The influence of track reconstruction on the collision vertex is examined by comparing the results for different $|z_{\mathrm{vtx}}|$ cuts, with differences found to be $0.5$–3% for all observables.
-Comparisons between different data-taking periods, particularly normal and reverse magnetic field runs in Au+Au collisions, show consistency within their statistical uncertainties.
+The influence of track reconstruction on the collision vertex is examined by comparing the results for different $|z_{\mathrm{vtx}}|$ cuts, with variations found to be $0.5$–3% for all observables.
+Comparisons between data-taking periods, particularly normal and reverse magnetic field runs in Au+Au collisions, show consistency within their statistical uncertainties.
 The influence of pileup and background events is studied by varying the cut on the correlation between $N_{\mathrm{ch}}^{\mathrm{rec}}$ and the number of hits in the TOF.
@@
 Subsequently, a track-by-track weight, as described in Eq.
-9, was applied to the accepted particles.
-All observables are calculated using the accepted particles and compared to those obtained using the original particles.
+(9), was applied to the accepted particles.
+All observables are calculated using the accepted particles and compared with those obtained using the original particles.
 This procedure allowed us to recover $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ within their statistical uncertainties.
 However, a 2–3% nonclosure was observed in $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$.
-Nevertheless, it’s important to note that such non-closures largely cancel out when considering the ratios between U+U and Au+Au collisions.
-Several additional cross-checks were also carried out.
+Nevertheless, it’s important to note that such non-closures largely cancel when considering the ratios between U+U and Au+Au collisions.
+Several additional cross-checks were carried out.
 The track reconstruction efficiency has an approximate 5% uncertainty due to its reliance on particle type and occupancy dependence.
-We repeated the analysis by varying this efficiency, and the variations in the results were either less than 1% or consistent with their statistical uncertainties.
+We repeated the analysis by varying this efficiency, and the variations in the results were either less than 1% or consistent within their statistical uncertainties.
 The reconstructed $p_{\mathrm{T}}$ can differ from the true value due to finite momentum resolution.
-This effect was investigated by further smearing the reconstructed $p_{\mathrm{T}}$ according to the known resolution, calculating the observable, and comparing the results with the original ones.
+This effect was investigated by smearing the reconstructed $p_{\mathrm{T}}$ according to the known resolution, calculating the observable, and comparing the results with the original ones.
 A discrepancy of approximately 0.5% was observed for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, while other observables remained consistent within their statistical uncertainties.
 These effects cancel in the ratios between U+U and Au+Au collisions.
-Modest differences are observed between the standard method and subvent methods for all observables, particularly towards the peripheral collisions.
-Some of these differences are attributable to non-flow contributions, while the remainder could arise from longitudinal flow decorrelations.
-We obtain default results as the average of the standard method and two-subevent method, and half of the differences are assigned as systematic uncertainties.
-They are 1–2% for $\left\langle v_{2}^{2}\right\rangle$, 1–3% for $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and 2–4% for $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in the 0–40% centrality range, respectively.
-The total systematic uncertainties are combined in quadrature from different sources, and in some cases, they are larger than the statistical uncertainties.
-The total uncertainties of $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are 2.5–4%, 2–5%, and 4–10%, respectively.
-Note that for the ratios between U+U and Au+Au, their uncertainties are evaluated for each source and combined in quadrature to form the total systematic uncertainties.
+The default results are obtained from the two-subevent method.
+The total systematic uncertainties, including these sources and non-flow, are calculated as a function of centrality.
+The uncertainties of the ratios between U+U and Au+Au are evaluated for each source and combined in quadrature to form the total systematic uncertainties.
 This process results in a partial cancellation of the uncertainties between the two systems.
-.9 Assigning uncertainties on $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$
+The uncertainties from different sources discussed above on the ratios are shown by the black boxes in Fig.
+6.
+The total systematic uncertainties including non-flow in the 0–5% centrality range amount to 3.9%, 4.4%, and 12.5% for $R_{v_{2}^{2}}$, $R_{(\delta p_{\mathrm{T}})^{2}}$, and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, respectively.
+.10 Hydrodynamic model setup and simulation
+Table 1 details the Woods-Saxon parameters for Au and U used in the IP-Glasma+MUSIC model calculations.
+The nucleon-nucleon inelastic cross-sections are the standard values 42 mb and 40.6 mb for Au+Au collisions at 200 GeV and U+U collisions at 193 GeV, respectively.
+For U, the nuclear shape in Eq.
+(1) is extended to include a possible small axial hexadecapole deformation $\beta_{4}$:
+$R(\theta,\phi)$ $=R_{0}(1+\beta_{2}[\cos\gamma Y_{2,0}+\sin\gamma Y_{2,2}]+\beta_{4}Y_{4,0})\;.$ (11)
+Most low-energy nuclear structure models favor a modest oblate deformation for 197Au Bally et al.
+(2023).
+We assume $\beta_{2\mathrm{Au}}=0.14$ and and $\gamma_{\mathrm{Au}}=45^{\circ}$ as the default choice for 197Au, which are varied within the range in the range of $\beta_{2\mathrm{Au}}\approx 0.12-0.14$ and $\gamma_{\mathrm{Au}}\approx 37-53^{\circ}$ according to Refs.
+Ryssens et al.
+(2023); Bally et al.
+(2023).
+These calculations reasonably reproduce many observables related to the ground-state nuclear deformation.
+For 238U, we scan several $\beta_{2\mathrm{U}}$ values ranging from 0 to 0.34.
+We also vary $\beta_{4\mathrm{U}}$ from 0 to 0.09 and $\gamma_{\mathrm{U}}$ in the range of 0–20 to examine the sensitivity of the U+U results to hexadecapole deformation and triaxiality.
+For each setting, about 100,000–400,000 events are generated using the officially available IP-Glasma+MUSIC Schenke et al.
+(2020a, b).
+Each event is oversampled at least 100 times to minimize statistical fluctuations in the hadronic transport.
+These calculations were performed using services provided by the Open Science Grid (OSG) Consortium Pordes et al.
+(2007); Sfiligoi et al.
+(2009).
+The role of final state effects is studied by varying the shear and bulk viscosities simultaneously up and down by 50%.
+The impacts on $\left\langle v_{2}^{2}\right\rangle$, $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$, and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are shown for Au+Au collisions in top of panels of Fig.
+7.
+The values of these flow observables are changed by more than a factor of two as a function of centrality.
+Yet, the ratios between U+U and Au+Au collisions, shown in the bottom panels, are relatively stable.
+A small reduction of $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ are observed in non-central collisions, when values of viscosities are halved.
+However, this change is an overestimate since the calculated flow observables greatly overestimate the data.
+So, in the end, half of the variations of the ratios are included in the model uncertainty.
+Fig.
+7: Sensitivity to shear and bulk viscosities of the QGP.
+IP-Glasma+MUSIC model prediction of the $\left\langle v_{2}^{2}\right\rangle$ (left), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (middle) and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (right) compared with the data in Au+Au collisions and the corresponding ratios (bottom) between U+U and Au+Au collisions, for different amount of shear and bulk viscosities.
+The main theoretical uncertainties arise from variations in nuclear structure parameters.
+Parameters common between two collision systems, such as the minimum inter-nucleon distance in nuclei $d_{\mathrm{min}}$, are not expected to contribute to the uncertainty significantly.
+However, other parameters, including nuclear radius $R_{0}$, skin $a$, and higher-order hexadecapole deformation $\beta_{4}$, could be different between Au and U and hence contribute more to the theoretical uncertainty.
+Table 1 provides a list of variations of nuclear structure parameters.
+The impact of these variations on ratios of flow observables is displayed in Fig.
+8.
+The ratios of flow observables are insensitive to these variations in the most central collisions.
+$R_{v_{2}^{2}}$ is particularly sensitive to skin parameter $a$.
+This is understandable, as $v_{2}$ has a large contribution from the reaction plane flow, which varies strongly with the value of $a$ Jia et al.
+(2023).
+Fig.
+8: Sensitivity to nuclear structure parameters.
+IP-Glasma+MUSIC model prediction of the ratios of $\left\langle v_{2}^{2}\right\rangle$ (a), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (b), and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (c) between U+U and Au+Au collisions.
+The calculations are done for different Glauber model parameters, and they are compared with the data.
+Fig.
+9: Comparison between two hydrodynamic models.
+The ratios of $\left\langle v_{2}^{2}\right\rangle$ (left), $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ (middle) and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ (right) as a function of centrality from IP-Glasma+MUSIC (solid lines) and Trajectum (dashed lines), assuming $\beta_{2\rm U}=0.28$ (red) and $\beta_{2\rm U}=0.25$ (blue).
+For the IP-Glasma+MUSIC model, only the uncertainties for the default configuration are shown for clarity.
+They are compared with the data.
+Model uncertainties for the ratios are derived by combining the impact of varying viscosities, together with various sources from Fig.
+8.
+As a consequence, checks that are consistent with the default calculation within their statistical uncertainties do not contribute to the model uncertainties.
+The combined model uncertainties for one standard deviation are displayed in Fig.
+3.
+A cross-check is conducted for an alternative hydrodynamic code, the Trajectum model Nijs and van der Schee (2023); Giacalone et al.
+(2023).
+This model has 20 parameter sets obtained from a Bayesian analysis of the Pb+Pb data at the LHC, but was not tuned to the RHIC data.
+For this calculation, we simply repeat the calculation at RHIC energy and calculate the same observables.
+While the description of the $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ is reasonable, several parameter sets give negative values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in mid-central collisions, and subsequently not used.
+The calculation is performed for the remaining 16 parameter sets as a function of centrality, and root mean square variations among these calculations are assigned as the uncertainty.
+Figure 9 displays the ratios of flow observables from Trajectum and compares them to IP-Glasma+MUSIC.
+The results from these two models agree within their uncertainties for $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$, with Trajectum predictions slightly higher in the UCC region.
+This leads to slightly lower values of $\beta_{2\mathrm{U}}$ than the IP-Glasma model: $\beta_{2\mathrm{U}}=0.228\pm 0.013$ for $R_{v_{2}^{2}}$ and $\beta_{2\mathrm{U}}=0.276\pm 0.018$ for $R_{(\delta p_{\mathrm{T}})^{2}}$.
+For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, however, the Trajectum model tends to systematically under-predict the data, as well as has much larger uncertainties compared with IP-Glasma model.
+In central collisions, such discrepancy can be improved by using a larger triaxiality parameter value $\gamma_{\mathrm{U}}\sim 15^{\circ}$.
+Overall, the comparison of the Trajectum model with data gives similar constraints on $\beta_{2\mathrm{U}}$ with comparable uncertainties but larger $\gamma_{\mathrm{U}}$ value with bigger uncertainties (next section).
+Table 1: Choices of Woods-Saxon parameters in Eqs.
+(3) and (11), including deformations, in the IP-Glasma+MUSIC model.
+The default values are denoted by bold font, and the rest are variations designed to constrain the values of $(\beta_{2\mathrm{U}},\gamma_{\mathrm{U}})$ and derive theoretical uncertainties associated with other structure parameters.
+Species R(fm) a (fm) $d_{\mathrm{min}}$(fm) $\beta_{2}$ $\beta_{4}$ $\gamma$ ()
+197Au 6.62,6.38 0.52 0.9, 0.4 0.14,0.12 0 53, 45, 37
+238U 6.81,7.07 0.55 0.9, 0.4 0,0.15 0.09,0 0, 10
+0.495 0.22, 0.25 15, 20
+0.605 0.28,0.34
+.11 Assigning uncertainties on $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$
 A standard pseudo-experiment procedure, similar to that described in Ref.
@@
 Here, $\bar{\beta}_{a}=0.294$ and $\sigma_{a}=0.021$ represent the mean and uncertainty of $\beta_{2\mathrm{U}}$ extracted from $R_{(\delta p_{\mathrm{T}})^{2}}$ in Fig.
-3g.
+3g from the IP-Glasma+MUSIC model.
 Similarly, $\bar{\beta}_{b}$ and $\sigma_{b}$ are the mean and uncertainty of $\beta_{2\mathrm{U}}$ from $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, and they depend on the parameter $\gamma_{\mathrm{U}}$.
 We sample a uniform prior distribution in $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$ to obtain the posterior distribution.
-From this posterior distribution, we obtained the mean and one standard deviation uncertainty of $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$, as well as the confidence contours displayed in Fig.
+From this posterior distribution, we obtained the mean and one standard deviation uncertainty of $\beta_{2\mathrm{U}}$ and $\gamma_{\mathrm{U}}$, $\beta_{2\mathrm{U}}=0.297\pm 0.015$ and $\gamma_{\mathrm{U}}=8.5^{\circ}\pm 4.8^{\circ}$, as well as the confidence contours displayed in Fig.
 3g.
-We also performed an extraction using all three ratio observables.
-In this case, an additional Gaussian probability density function derived for $R_{v_{2}^{2}}$ with a mean $\bar{\beta}_{c}=0.240$ and variance $\sigma_{c}=0.014$ is multiplied to Eq.
-(10), and the same procedure described above is repeated.
-This gives $\beta_{2\mathrm{U}}=0.281\pm 0.010$ and $\gamma_{\mathrm{U}}=4.4^{\circ}\pm 3.2^{\circ}$.
-However, we do not use this estimate, given the imperfect description of $v_{2}$ in the hydrodynamic model.
-.10 Hydrodynamic model setup and simulation
-Table 1 details the Woods-Saxon parameters for Au and U used in the IP-Glasma+MUSIC model calculations.
-The nucleon-nucleon inelastic cross-sections are chosen to be the standard values 42 mb and 40.6 mb for Au+Au collisions at 200 GeV and U+U collisions at 193 GeV, respectively.
-For U, the nuclear shape in Eq.
-(1) is extended to include a possible small axial hexadecapole deformation $\beta_{4}$:
-$R(\theta,\phi)$ $=R_{0}(1+\beta_{2}[\cos\gamma Y_{2,0}+\sin\gamma Y_{2,2}]+\beta_{4}Y_{4,0})\;.$ (11)
-Most low-energy nuclear structure models favor a modest oblate deformation for ${}^{197}$u Bally et al.
-(2023) in the range of $\beta_{2\mathrm{Au}}\approx 0.12-0.14$.
-We assume ${}^{197}$u is oblate shape with $\beta_{2\mathrm{Au}}=0.14$ as the default choice for simulation.
-For ${}^{238}$, we have scanned several $\beta_{2\mathrm{U}}$ values ranging from 0 to 0.34.
-We also vary $\beta_{4\mathrm{U}}$ from 0 to 0.09 and $\gamma_{\mathrm{U}}$ in the range of ${}^{\circ}$${}^{\circ}$ to examine the sensitivity of the U+U results to hexadecapole deformation and triaxiality.
-For each setting, about 100,000–400,000 events are generated using the officially available IP-Glasma+MUSIC Schenke et al.
-(2020a, b).
-Each event is oversampled at least 100 times to minimize non-flow effects and statistical fluctuations in the hadronic transport.
-These calculations were performed using services provided by the Open Science Grid (OSG) Consortium Pordes et al.
-(2007); Sfiligoi et al.
-(2009).
-The role of final state effects is studied by varying the shear and bulk viscosities simultaneously up and down by 50%.
-The impact on $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ and $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ are shown for Au+Au collisions in top of panels of Fig.
-6.
-The values of these flow observables are changed by more than a factor of two as a function of centrality, yet the ratios between U+U and Au+Au collisions, shown in the bottom panels, are relatively stable.
-A small reduction of $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$ are observed in the non-central collisions, when values of viscosities are halved.
-However, this change is clearly an overestimate since the calculated flow observables greatly overestimate the data.
-So in the end, half of the variations of the ratios are included in the model uncertainty.
-The main theoretical uncertainties arise from variations in nuclear structure parameters.
-Parameters common between two collision systems, such as the minimum inter-nucleon distance in nuclei $d_{\mathrm{min}}$, are not expected to contribute to the uncertainty significantly.
-However, other parameters including nuclear radius $R_{0}$, skin $a$, and higher-order deformation $\beta_{4}$ could be different between Au and U, and hence contribute more to the theoretical uncertainty.
-Table 1 provides a list of variations of nuclear structure parameters.
-The impact of these variations on ratios of flow observables is displayed in Fig.
-7.
-The ratios of flow observables are insensitive to these variations in the most central collisions.
-Clearly, $R_{v_{2}^{2}}$ is particularly sensitive to skin parameter $a$.
-This is understandable, as $v_{2}$ has a large contribution from the reaction plane flow, which varies strongly with the value of $a$ Jia et al.
-(2023).
-Model uncertainties for the ratios are derived by combining the impact of varying viscosities, together with various sources from Fig.
-7 according to the Barlow prescription Barlow (2002).
-As a consequence, checks that are consistent with the default calculation within their statistical uncertainties do not contribute to the model uncertainties.
-The combined model uncertainties for one standard deviation are illustrated in Fig.
-3.
-A cross-check was also carried out for an alternative hydrodynamic code, the Trajectum model Nijs and van der Schee (2023); Giacalone et al.
-(2023).
-This model has 20 parameter sets obtained from a Bayesian analysis of the Pb+Pb data at the LHC, but was unfortunately not tuned to the RHIC data.
-For this calculation, we simply repeat the calculation at RHIC energy and calculate the same observables.
-While the description of the $\left\langle v_{2}^{2}\right\rangle$ and $\left\langle(\delta p_{\mathrm{T}})^{2}\right\rangle$ is reasonable, several parameter sets give negative values of $\left\langle v_{2}^{2}\delta p_{\mathrm{T}}\right\rangle$ in mid-central collisions, and subsequently not used.
-The calculation is performed for the remaining 16 parameter sets as a function of centrality, and root mean square variation among these calculations is assigned as the uncertainty.
-Figure 8 displays the ratios of flow observables from Trajectum and compares them to IP-Glasma+MUSIC.
-The results from these two models agree within their uncertainties for $R_{v_{2}^{2}}$ and $R_{(\delta p_{\mathrm{T}})^{2}}$, with Trajectum predictions slightly higher in the UCC region.
-This leads to slightly lower values of $\beta_{2\mathrm{U}}$: $\beta_{2\mathrm{U}}=0.230\pm 0.010$ for $R_{v_{2}^{2}}$ and $\beta_{2\mathrm{U}}=0.273\pm 0.015$ for $R_{(\delta p_{\mathrm{T}})^{2}}$.
-For $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$, however, the Trajectum model tends to under-predict the data.
-In central collisions, such discrepancy can be improved by adjusting the triaxiality parameter to $\gamma_{\mathrm{U}}=15^{\circ}$.
-Overall, the comparison of the Trajectum model with data leads to similar constraints on $\beta_{2\mathrm{U}}$ but slightly larger $\gamma_{\mathrm{U}}$.
-Table 1: The choices of Woods-Saxon parameters in Eqs.
-(3) and (11), including deformations, in the IP-Glasma+MUSIC model.
-The default values are denoted by bold font, while the rest are variations designed to constrain the values of $(\beta_{2\mathrm{U}},\gamma_{\mathrm{U}})$ and derive theoretical uncertainties associated with other structure parameters.
-Species R(fm) a (fm) $d_{\mathrm{min}}$(fm) $\beta_{2}$ $\beta_{4}$ $\gamma$ (${}^{\circ}$)
-${}^{197}$u 6.62,6.38 0.52 0.9, 0.4 0.14,0.12 0 60
-${}^{238}$ 6.81,7.07 0.55 0.9, 0.4 0,0.15,0.22 0.09,0 0, 10, 15
-0.495,0.605 0.25,0.28,0.34 20
+This statistical analysis is also performed for $R_{(\delta p_{\mathrm{T}})^{2}}$ and $R_{v_{2}^{2}\delta p_{\mathrm{T}}}$ for the Trajectum model, yielding constraints of $\beta_{2\mathrm{U}}=0.275\pm 0.017$ and $\gamma_{\mathrm{U}}=15.5^{\circ}\pm 7.8^{\circ}$.
+Finally, we perform an analysis to combine the constraints of the IP-Glasma+MUSIC and Trajectum models.
+This is achieved by multiplying the probability density function Eq.
+(12) from the two models, treating their constraints as statistically independent.
+This approach yields $\beta_{2\mathrm{U}}=0.286\pm 0.012$ and $\gamma_{\mathrm{U}}=8.7^{\circ}\pm 4.5^{\circ}$.
+We noticed that Trajectum model does not impact the constraints on $\gamma_{\mathrm{U}}$ due to the model’s large uncertainty, but the uncertainty on $\beta_{2\mathrm{U}}$ reduces significantly due to the comparable precision in the two models.
+Therefore, we also include the difference of the extracted $\beta_{2\mathrm{U}}$ values between the two models as an additional theoretical uncertainty.
+The final constraints given by this procedure are $\beta_{2\mathrm{U}}=0.286\pm 0.025$ and $\gamma_{\mathrm{U}}=8.7^{\circ}\pm 4.5^{\circ}$.
+.12 Data availability
+All raw data for this study were collected using the STAR detector at Brookhaven National Laboratory and are not available to the public.
+Derived data supporting the findings of this study are publicly available in the HEPData repository (https://www.hepdata.net/record/147196) or from the corresponding author on request.
+.13 Code availability
+The codes to process raw data collected by the STAR detector are publicly available on GitHub (https://github.com/star-bnl). The codes to analyse the produced data are not publicly available.